### v1

# Incorporating Judgment Prediction into Legal Case Retrieval via Law-aware Generative Retrieval

Abstract.
Legal case retrieval and judgment prediction are crucial components in intelligent legal systems. In practice, determining whether two cases share the same charges through legal judgment prediction is essential for establishing their relevance in case retrieval. However, current studies on legal case retrieval merely focus on the semantic similarity between paired cases, ignoring their charge-level consistency. This separation leads to a lack of context and potential inaccuracies in the case retrieval that can undermine trust in the system’s decision-making process. Given the guidance role of laws to both tasks and inspired by the success of generative retrieval, in this work, we propose to incorporate judgment prediction into legal case retrieval, achieving a novel law-aware Generative legal case retrieval method called Gear. Specifically, Gear first extracts rationales (key circumstances and key elements) for legal cases according to the definition of charges in laws, ensuring a shared and informative representation for both tasks. Then in accordance with the inherent hierarchy of laws, we construct a law structure constraint tree and assign law-aware semantic identifier(s) to each case based on this tree. These designs enable a unified traversal from the root, through intermediate charge nodes, to case-specific leaf nodes, which respectively correspond to two tasks. Additionally, in the training, we also introduce a revision loss that jointly minimizes the discrepancy between the identifiers of predicted and labeled charges as well as retrieved cases, improving the accuracy and consistency for both tasks. Extensive experiments on two datasets demonstrate that Gear consistently outperforms state-of-the-art methods in legal case retrieval while maintaining competitive judgment prediction performance.
Legal case retrieval, Generative Retrieval, Legal Judgment Prediction
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Legal case retrieval, which seeks to retrieve relevant precedent cases based on a usually undecided case, stands as a pivotal component in intelligent legal systems. This task relies on specialized legal knowledge encoded in laws, with the goal of providing essential evidence to legal practitioners and supporting their decision-making processes.
Existing studies on legal case retrieval (Ma et al., 2021; Xiao et al., 2021a; Chalkidis et al., 2020) have been focused on the semantic similarity between paired cases, for example,  Li et al. (2023) utilize the structural information contained in legal cases to conduct fine-grained semantic interactions between the query and candidate cases. Although they have achieved impressive performances, they ignore the fact that legal case retrieval and judgment prediction [^1] Legal judgment prediction aims to predict the relevant law article, charge, and term of penalty for a given query case. In this work, we focus on the charge prediction. Other sub-tasks are easily incorporated into our proposed method. are closely intertwined (Shao et al., 2023; Zhang and Dou, 2023). In particular, the charge-level consistency between the query and retrieved case, a direct product of the judgment prediction, is crucial for relevance estimation in legal case retrieval. As shown in Figure 1, it provides valuable contextual information and effective relevance signals. Moreover, incorporating judgment prediction into case retrieval improves the transparency (Olteanu et al., 2021; Kim et al., 2020; Baan et al., 2019) of the decision-making process. This integration aids legal practitioners in tracing the retrieved cases from the perspective of charges, thereby making the legal decisions trustworthy.
Figure 1. A pair of relevant cases from the LeCaRDv2 dataset. The applicable charges (underlined) and their corresponding rationales (red and blue) provide effective relevance signals.
Despite the aforementioned benefits, there remain the following challenges in incorporating judgment prediction into case retrieval. Firstly, legal case retrieval and judgment prediction are usually formulated as two distinct machine learning problems–ranking and classification. It is difficult for one ranking model to predict the applicable charges for legal cases and leverage the charge prediction to enhance ranking. Secondly, legal cases are inherently lengthy and complicated. As shown in Figure 1, only the rationales (key circumstances and key elements)  (Ma et al., 2021; Yu et al., 2022) of the case provide effective signals for relevance estimation. It is expected to represent each case as a shared and informative representation for both tasks, but existing methods often encode legal cases in a coarse-grained granularity way (Chalkidis et al., 2020; Xiao et al., 2021a). Thirdly, both tasks rely on specialized law knowledge (Sun et al., 2022; Liu et al., 2023c), an appropriate way that effectively incorporates law expertise to guide the prediction of both tasks remains a concern, given that current researches (Chalkidis et al., 2020; Xiao et al., 2021a; Zhang et al., 2023a; Li et al., 2023) solely focus on text-level semantics for predictions.
Facing the above challenges and inspired by the success of generative retrieval (Tay et al., 2022; Wang et al., 2022), we propose a novel generative legal case retrieval method, namely Gear, to incorporate judgment prediction into legal case retrieval. As illustrated in Figure 2, Gear unifies these two tasks into a tree-based generative retrieval framework. The insight of Gear is to map the query case to a set of law-aware semantic identifiers where each identifier represents a precedent case relevant to the query and also reflects applicable charges. Specifically, we first construct a corpus based on the applicable laws [^2] We consider the Criminal Law of the People’s Republic of China in this work. and subsequently extract rationales at word- and sentence-level for each case according to this corpus. It effectively filters out the noise and ensures the case representations are shared and informative to both tasks. Then, we create the law structure constraint tree based on the inherent hierarchy of laws (Chapter-Section-Article) , considering that both tasks are learned with the guidance of laws. Thus, we can consider situations where each case may involve single or multiple charges and assign the law-aware semantic identifier(s) to each case given this tree, making the identifier reflect its charges. In this way, the retrieval is equivalent to a traversal from the root, through intermediate charge nodes, to case-specific leaf nodes, which respectively correspond to two tasks. To improve the accuracy and consistency for both tasks, we develop a revision loss that jointly minimizes the discrepancy between predicted and labeled charges/ retrieved cases. Furthermore, to assess the performance of judgment prediction in case retrieval, we introduce a novel metric called coverage. This metric gauges the charge-level consistency between the query and retrieved cases.
Figure 2. The overview of our proposed Gear. It mainly consists of three modules, including rationale extraction, law structure constraint tree, and revision loss. The middle part in blue is the generative retrieval framework.
The major contributions of the paper are summarized as follows:
(1) To the best of our knowledge, this is the first work that incorporates judgment prediction into legal case retrieval, addressing the challenge of charge-level consistency between the query case and retrieved cases in legal search. Our work also improves the transparency of the legal decision-making process.
(2) We propose a novel model, namely Gear, to unify legal case retrieval and judgment prediction under the generative retrieval framework. We explicitly leverage the law knowledge to extract rationales from cases, assign cases the law-aware hierarchical identifies, and formulate the prediction as a traversal on the law structure constraint tree. We also propose a revision loss to jointly improve the accuracy and consistency of both tasks.
(3) We conduct extensive experiments on two public datasets and propose an evaluation metric for assessing the charge-level consistency between the query and retrieved cases. Experimental results indicate Gear not only achieves state-of-the-art performance in legal case retrieval but also maintains competitive judgment prediction performance.
2. Methodology
2.1. Task Formulation
In this work, we target on the task of legal case retrieval. Suppose that we have a set of collected samples $\mathcal{D}=\{(q,\mathcal{C},\mathcal{R})\}$. For each data instance, $q$ represents a undecided query case submitted by the legal practitioner; $\mathcal{C}=\{c_{1},c_{2},\cdots,c_{N}\}$ represents candidate cases (precedents) pool with size $N\in\mathbb{N}^{+}$; $\mathcal{R}$ represents the labeled relevant case set from $\mathcal{C}$ given the query case $q$. Unlike previous studies that only predict to retrieve $\mathcal{R}$ from $\mathcal{C}$ given $q$, in this work, we instead unify judgment prediction and case retrieval into a generative retrieval framework, and thus aims at learning a ranking function $f:q\times\mathcal{C}\rightarrow\mathcal{R}\times\mathcal{E}$, where $\mathcal{E}$ denotes the set of applicable charges corresponding to $q$.
2.2. The Principle of Our Method
To retrieve relevant cases and predict their applicable charges, we develop a novel learning method from the viewpoint of generative retrieval (GR). Essentially, GR is an emerging new retrieval paradigm, that adopts a deep neural network to perform the end-to-end retrieval in contrast to traditional the index-then-retrieve paradigm. GR models consist of two major steps to directly generate identifier strings of documents as the retrieval target. In the first indexing step that focuses on memorizing the information about each document, GR models take each original document $c$ as input and generate its identifier ${id}^{c}$ as output in a straightforward Seq2Seq fashion. The model is trained with the standard language model objective with the teacher forcing:
(1) $\mathcal{L}_{i}=\sum_{c\in\mathcal{C}}\log P({id}^{c}|c),$
where ${id}^{c}$ is represented by the atomic identifier or the semantic identifier. Atomic identifiers are generated by assigning each document an arbitrary integer, while semantic identifiers are generated according to the hierarchical clustering algorithm employed over all the documents. In the second retrieval phase, GR models associate each $q$ to its relevant document $r$ through an auto-regressive generation:
(2) $\mathcal{L}_{r}=\sum_{(q,r)\in\mathcal{D}}\log P({id}^{r}|q),$
where ${id}^{r}$ denotes the identifier of $r$. ${id}^{r}$ is also represented by the atomic identifier or the semantic identifier. As such, once a GR model is learned, it can be used to retrieve candidate documents for a test query in an end-to-end manner using beam search.
As aforementioned in Section 1, applying GR to our legal scenario faces several challenges including the lengthy and complicated content, the law-aware semantic identifier, and the training objective for enhancing both tasks. Therefore, when incorporating legal judgment prediction into case retrieval in the above GR framework, the critical learning tasks become: (1) extract rationales instead of using original documents to form the input for GR (Section 2.3); (2) create informative law-aware identifiers for each case based on the hierarchical structure of laws (Section 2.4); (3) develop a training objective to explicitly ensure charge-level and case-level consistency between predictions and labels(Section 2.5).
Figure 3. The proposed three modules of Gear. $f_{R}$ extract rationales from legal cases; $f_{T}$ assigns hierarchical identifiers to each case and constrain the decoding; $\mathcal{L}_{c}$ jointly optimizes case retrieval and judgement prediction.
2.3. Rationale Extraction
To filter the noise and provide shared and informative representations for legal cases, based on the set of laws $L$, we devise a module called $f_{R}$ to extract rationales $E$ instead of using the original document $doc\in\{q,c\}$ as the input of Gear:
(3) $E=f_{R}(doc,L),$
where $E=\{E_{w},E_{s}\}$ in which $E_{w}$ and $E_{s}$ respectively denotes the word-level and sentence-level rationales.
As shown in Figure 3, to achieve this goal, we first leverage the guidance of laws to construct a corpus $B$, collecting law-based keywords and embeddings. Specifically, the corpus consists of three parts $B=\{B_{1},B_{2},B_{3}\}$. $B_{1}$ is constructed according to the lexical variants of all charges names in $L$, for example, in terms of the charge “crime of forges the seals of a company, enterprise, institution or a people’s organization”(translated from Chinese), we split the charge name and put all keywords and their lexical variants to the corpus after the stop words are removed. $B_{2}$ is constructed similarly to the $B_{1}$. We put all keywords in the definition of laws and the feedback from prompting ChatGPT [^3] https://openai.com/chatgpt, the prompts are provided in Appendix B.2. to the corpus after the stop words are removed. $B_{3}$ is constructed by collecting BERT (Devlin et al., 2018; Chalkidis et al., 2020) embeddings [^4] https://github.com/thunlp/OpenCLaP. for the definition of each law.
Once the corpus is collected, we compute the multiple-level scores and extract $E$ for each case based on three corpora as follows. At the word level, we split the case and respectively select the top-$k_{1}$ (top-$k_{2}$) keywords $E_{w_{1}}$ ($E_{w_{2}}$) from $B_{1}$ ($B_{2}$) as follows:
(4) $E_{w_{i}}=\mathop{\arg\text{top-}k_{i}}_{w\in B_{i}}\biggl{(}\text{tf}(w,doc)\biggr{)},$
where $i\in\{1,2\}$; $w$ is the word from $B_{i}$; $k_{1}$ and $k_{2}$ are hyperparameters to control the number of selected words; tf($\cdot$) denotes term frequency. At the sentence level, for each sentence $s$ in $doc$, we first select $E_{s}$ from as follows:
(5) $E_{s}=\mathop{\arg\text{top-}k_{3}}_{s\in doc,l\in L}\biggl{(}$ $\lambda_{1}\frac{\text{sim}(s,B_{1})}{\text{len}(s)}+\lambda_{2}\frac{\text{sim}(s,B_{2})}{\text{len}(s)}+$
$\lambda_{3}\cos(\text{emb}(s),\text{emb}(l))\biggr{)},$
where $k_{3}$ is another hyperparameter to control the number of selected sentences; $\lambda_{1},\lambda_{2},\lambda_{3}$ are balance coefficients; len() denotes the sentence length; $\cos(\cdot)$ denotes the cosine similarity; emb() denotes the embedding function; sim($\cdot$) is defined as:
(6) $\text{sim}(s,B_{i})=\sum_{w\in s}\text{tf}(w,B_{i}),$
where $w$ is the word from $s$. Please note that since query cases are typically undecided with no charges, we extract rationales using the method described above. On the other hand, candidate cases are precedents with decided charges, thus, when extracting the rationales for them, we shrunk $B$ into a corpus constructed based on its corresponding charges.
2.4. Law Structure Constraint Tree
Considering that both legal case retrieval and judgment prediction require guidance by the law where there usually exists a “Chapter-Section-Article” hierarchy, we argue that the decision process to judge whether two cases are relevant in case retrieval is analogous to the search in such tree-like hierarchy. In other words, when the practitioners search for the relevant cases given a query case, they always expect the charge of the relevant precedent case to be located at the same position in the hierarchy as the charges applicable in the query case. Therefore, we devise a module $f_{T}$ that leverages the inherent hierarchy of the law to construct a law structure constraint tree $T$ as shown in Figure 3 and assigns the law-aware semantic identifier $id$ for each $c\in\mathcal{C}$:
(7) $id=f_{T}(c,T),$
where $id$ is in the prefix-suffix style. The prefix depends on the position of the applicable charges within the tree. As illustrated in Figure 3, case 809 involves the charge of “crime of robbery” which falls under Article 269 of Chapter 5 of Section 2 of the Criminal Law of the People’s Republic of China. Hence, the assigned prefix for this case is 0-2-5-269, where 0 represents the root node of the tree. Regarding the suffix, we regard legal cases are the children of their corresponding charges on the tree, and assign a unique identifier to each case under a crime node. Compared to current works that employ hierarchical $k$-means to create identifiers for each document, ours avoid that same integers may have different meanings at different levels, and thus ensure the effectiveness of model learning.
In the retrieval (inference), the identifiers of top relevant cases can be easily generated via beam search. Due to the hierarchical nature of laws as analyzed above, it is convincing to use constrained beam search (Anderson et al., 2017; Hu et al., 2019; Post and Vilar, 2018) decoding process with our law structure constraint tree, which in turn only generates the valid identifiers. On the other hand, one unique feature of the legal domain is that a single case can involve multiple charges. When the query is this kind of case, the ideal retrieval results should encompass these charges. Therefore, we assign $k$ identifiers to cases involving $k$ charges, with each identifier corresponding to a specific charge. For example, as shown in Figure 3, the case 725 involves charge 267 and charge 269, the valid identifiers for this case include 0-2-5-267-725 and 0-2-5-269-725. In doing so, during the retrieval, we expect the model to retrieve all identifiers of the target case, thereby increasing the probability of the target being retrieved.
2.5. Revision Loss
Besides employing the typical language model training objectives (Equation 1 and Equation 2), we also develop a revision loss for the consistency between the query and retrieved cases, i.e., we aim at directly minimizing their charge-level and case-level discrepancy.
Formally, as illustrated in Figure 3, given the predicted identifier (list of integers) $[\hat{id}_{1},\hat{id}_{2},\cdots,\hat{id}_{L}]$ and the corresponding ground-truth identifier $[{id}_{1},{id}_{2},\cdots,{id}_{L}]$ for a query $q$ at the same position of the ranking, both having a length of $L$, we compare the difference between them and calculate the reward $R_{t}$ at each step $t\in[1,L]$ as follows:
(8) $R_{t}=\begin{cases}\mu,&\text{if}~{}\hat{id}t={id}_{t},\\ -\gamma_{h}^{L-t}\mu|\hat{id}_{t}-{id}_{t}|,&\text{if}~{}\hat{id}_{t}\neq{id}_{t},\\ \end{cases}$
where $\mu$ is the constant reward unit, $\gamma_{h}\in(0,1]$ is the hierarchy penalty factor used to penalize the differences between predictions and labels layer by layer along the law structure constraint tree, with larger penalties as it get closer to the tree root and smaller penalties as it get closer to the tree leaves. Intuitively, if two cases share the same prefix, they are likely relevant to each other because of the same applicable charge, thus receiving higher rewards. Conversely, if two cases belong to different chapters in terms of applicable charges, they are likely dissimilar resulting in lower rewards.
Then we apply the REINFORCE algorithm (Williams, 1992) to optimize the model parameters, the revision loss is defined to minimize the policy gradient objective:
(9) $\mathcal{L}_{c}=-\sum_{q\in\mathcal{D}}\sum^{bz}_{b}\gamma_{r}^{b}\sum^{L}_{t}\log p(\hat{id}_{t}|q)\cdot R_{t}-\lambda_{w}\log p({id}_{t}|q),$
where $\gamma_{r}\in(0,1]$ the optional rank penalty factor used to focus on the top of ranking; $bz$ denotes the beam size, i.e., the number of cases to retrieve for each query; $p(\hat{id}_{t}|q)$ is the probability that predicting to generate $\hat{id}_{t}$ given $q$ at layer $t$; $\lambda_{w}$ is the balance coefficient. To further handling the sparse reward issue and improve the training efficiency, we follow (Xin et al., 2020; Chen et al., 2021; Liu et al., 2023a) and add the second term to Equation 9 that directly increases the probability of generating ${id}_{t}$. Therefore, the overall training objective for Gear is:
(10) $\mathcal{L}=\mathcal{L}_{i}+\mathcal{L}_{r}+\lambda_{l}\mathcal{L}_{c},$
where $\lambda_{l}$ is the coefficient to balance the indexing loss (Equation 1), retrieval loss (Equation 2), and the revision loss.
2.6. Discussion
In this part, we highlight the merits and limitations of the proposed Gear approach as follows:
Transparency. Inspired by the decision-making process in legal scenarios, our approach formulates the legal case retrieval as the traversal on the law structure constraint tree. This design not only provides valuable contextual information but also improves the transparency of retrieval results by tracing back the decision-making process.
Versatility. Since the retrieved cases are represented by the law-aware semantic identifier with each digit denoting one prediction at each layer in the law structure constraint tree, we can directly predict the applicable judgments for the query case according to the retrieval results. This design makes it versatile to unify legal case retrieval and judgment prediction.
Limitation. In this work, we mainly focus on the Criminal Law of the People’s Republic of China. It follows the civil law system and has an explicit hierarchy. In countries that follow the common law system such as the United States and the United Kingdom, their laws usually do not possess an explicit hierarchical structure. It may limit the applicability of our Gear. Fortunately, for example, in the United States, there is a combination of federal and state laws, where the latter often follows a structure involving “titles, chapters, sub-chapters” hierarchy. Therefore, our method can be applied to such scenarios.
3. Experiments
In this section, we conduct experiments to answer the following research questions:
  • RQ1: How does Gear perform on legal case retrieval compared to state-of-the-art methods?
  • RQ2: How effective are the three modules in Gear?
  • RQ3: Can Gear show competitive performance on judgment prediction?
  • RQ4: What is the quality of the rationales extracted by Gear including effectiveness and efficiency?
  • RQ5: Can Gear incur less time overhead in legal case retrieval compared to popular generative methods?
3.1. Experimental Settings
3.1.1. Datasets
We conduct experiments on LeCaRDv2(Ma et al., 2021) and ELAM(Yu et al., 2022) for legal case retrieval. Table 1 reports basic statistics of both datasets. We leave the detailed descriptions in Appendix A.1.
Table 1. Statistics of ELAM and LeCaRDv2. “(Top 1)” indicates the charges corresponding to the most relevant candidate case serve as the label charges for the query case.
Dataset
Statistics ELAM LeCaRDv2
Avg. length per candidate case 1163.68 1568.38
Avg. length per query case 1304.98 558.18
Avg. # charges per candidate case 1.00 1.50
Avg. # charges per query case (Top 1) 1.00 1.53
# Train candidate cases/query 1332 1390
# Query cases involved 147 653
# Crimes involved 97 100
[…]
ELAM LeCaRDv2
Models R@1 R@5 R@10 R@20 MRR R@1 R@5 R@10 R@20 MRR
QL (Zhai et al., 2008) 0.0272 0.1088 0.1361 0.2857 0.0723 0.0252 0.0892 0.1351 0.2177 0.1327
BM25 (Robertson et al., 2009) 0.0340 0.0680 0.1497 0.2245 0.0635 0.0435 0.1452 0.2549 0.3900 0.1862
ANCE (Xiong et al., 2020) 0.0769 0.1894 0.2357 0.3300 0.1512 0.0362 0.0929 0.0968 0.1145 0.1323
BERT (Devlin et al., 2018)+ANN 0.0302 0.0908 0.1405 0.2461 0.0921 0.0299 0.0718 0.0757 0.0934 0.1162
Lawformer (Xiao et al., 2021a)+ANN 0.0484 0.1638 0.2209 0.3185 0.1410 0.0318 0.0591 0.0728 0.1093 0.1238
SAILER (Li et al., 2023) 0.0670 0.2132 0.3282 0.4604 0.1729 0.0208 0.0844 0.1510 0.2371 0.1318
DSI (Tay et al., 2022) 0.0204 0.1134 0.2274 0.3159 0.1249 0.0232 0.0577 0.0768 0.1285 0.1159
DSI-QG (Zhuang et al., 2022) 0.0278 0.1606 0.2736 0.3803 0.1531 0.0283 0.0725 0.1230 0.1881 0.1224
NCI (Wang et al., 2022) 0.0325 0.0936 0.1463 0.2070 0.1285 0.0416 0.1024 0.1696 0.2504 0.1914
Ultron (Zhou et al., 2022) 0.0607 0.1583 0.2260 0.3506 0.1678 0.0333 0.1207 0.2142 0.3492 0.1511
Gear ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
 w/o $\mathcal{L}_{c}$ 0.0611 0.1657 0.2793 0.4167 0.1763 0.0452 0.1485 0.2549 0.4478 0.1979
 w/o $f_{R}$ 0.0586 0.1830 0.2913 0.4437 0.1802 0.0308 0.1143 0.2108 0.3714 0.1621
 w/o $f_{T}$ 0.0464 0.1429 0.2328 0.3158 0.1626 0.0166 0.0550 0.0971 0.1388 0.1596
3.1.2. Baselines
We compare Gear with three types of representative baselines in this work including term-based methods, dense retrieval methods, and generative retrieval methods. We leave the introduction on baselines and the detailed implementation in Appendix A.2 and Appendix A.3, respectively.
3.1.3. Evaluation Metrics
Following previous works (Tay et al., 2022; Zhuang et al., 2022; Wang et al., 2022; Zhou et al., 2022), we leverage the commonly adopted metrics, including Recall (R) and MRR. The averaged results on all test cases are reported.
To demonstrate Gear’s ability on judgment prediction, we introduce a new metric called $coverage@k$, assessing the percentage of charges involved in the query cases that are covered by top-$k$ retrieved cases. This metric evaluates the charge-level consistency between the query case and the retrieved cases. Considering that query cases are usually undecided and LeCaRDv2 does not have the ground-truth judged charges for query cases, we adopt $coverage_{\text{freq}}$ and $coverage_{\text{top1}}$ to create a pseudo label for each query case. In terms of $coverage_{\text{freq}}$, for each query case, we create a charge dictionary from all the charges involved in all relevant candidate cases and select the charge that appears most frequently as the charge label for the query. In terms of $coverage_{\text{top1}}$, for each query case, we take the charge involved in the most relevant candidate case as the charge label for the query. In our experiments, we compute $coverage@k$ metric at 1, 3, 5, and 10.
3.2. Retrieval Performance
Table 2 presents the retrieval performance of Gear and baselines on the ELAM and LeCaRDv2. All the methods are trained 10 times and the averaged results are reported. From the results, we have the following observations for RQ1:
$\bullet$ Gear demonstrates a significant performance advantage over all baseline methods on both datasets. The relative improvements in terms of MRR on the ELAM and LeCaRDv2 are at least 36.77% and 16.09%, respectively. These results indicate Gear’s effectiveness. We contribute the improvement to the elaborate design for incorporating judgment into case retrieval including rationale extraction, identifier construction, and the revision loss. See Section 3.3 for detailed analysis for each module of Gear.
$\bullet$ Compared to sparse retrieval and dense retrieval methods, current generative retrieval methods struggle to achieve satisfying performance in the legal domain. Generative retrieval baselines merely focus on mapping original cases to their $k$-means style identifiers. Without injecting the law knowledge, these methods inevitably yield sub-optimal results. SAILER uses the case-level structure and obtains the second-best performance in ELAM where both query and candidate cases have clear structure. In LeCaRDv2 where only the fact part is provided for query case and there exhibits a distribution shift between query and candidate, BM25 performs second best. In contrast, Gear explicitly leverages the law knowledge to extract rationales, construct identifiers, and involve judgment prediction for legal case retrieval. These designs lead to significant performance improvement.
3.3. Ablation Studies
Figure 4. Comparison of retrieval performance of DSI and DSI equipped with the proposed three modules. The means values of 5 repeated experiments are reported, with error bars representing the 95% confidence interval of the means.
To answer RQ2, we conduct an ablation study to investigate the impact of each component of Gear.
Firstly, we test the performance of Gear’s variants by removing a certain component. Following DSI (Tay et al., 2022), we replace rationale extraction and law-aware hierarchical identifiers to the Direct indexing (the first 512 tokens) and $k$-means identifiers. As shown in the bottom part of Table 2, we observe that: (1) Removing modules $\mathcal{L}_{c}$, $f_{R}$, and $f_{T}$ individually results in a performance degradation of 25.51%, 23.83%, and 31.16% on ELAM, 8.50%, 25.00%, and 26.13% on LeCaRDv2. These results verify the effectiveness of all three modules. (2) It is worth noting that removing our law-aware hierarchical identifiers results in the most significant performance decrease. It is because we inject law knowledge into each digit of the identifier, aligning them with the hierarchical structure of laws. By simulating the retrieval to mirror the legal decision-making process, we enable Gear to effectively learn the association between query cases and candidate cases. These results demonstrate the effectiveness of our identifiers and highlight the importance of introducing the law knowledge to legal case retrieval.
To further validate the efficacy of the components of Gear, we combine the three components of Gear with DSI individually and test the retrieval performance. We conduct 5 repeated experiments and present the mean values of MRR, with error bars representing the 95% confidence interval. As illustrated in Figure 4, incorporating our components of Gear into DSI consistently results in performance improvements. Specifically, DSI equipped with rationale extraction achieves an improvement of over 60% on ELAM, and DSI equipped with law-aware identifiers demonstrated a 42% improvement on LeCaRDv2 dataset. However, DSI equipped with revision loss does not exhibit a substantial improvement. Because DSI utilizes $k$-means identifiers and the length of these identifiers may vary due to differences in cluster sizes. In such cases, our revision loss cannot accurately measure the differences between predictions and labels, resulting in a slight performance improvement.
(a) $coverage_{\text{freq}}$ on LeCaRDv2
(b) $coverage_{\text{top1}}$ on LeCaRDv2
(c) $coverage$ on ELAM
Figure 5. Comparison of the judgment prediction performances in terms of coverage. The proposed Gear consistently outperforms DSI and DSI-QG in both single (ELAM) and multiple (LeCaRDv2) charges scenarios.
3.4. Quality of Judgment Prediction
To answer RQ3, we conduct experiments to evaluate the judgment prediction accuracy of retrieved cases using the proposed metric $coverage$. Considering that in LeCaRDv2, the ground-truth judgment label of query cases is not provided, we use $coverage_{\text{freq}}$ and $coverage_{\text{top1}}$ to create pseudo labels for them. We run experiments ten times and present the results in Figure 5, reporting the average score with the shaded area indicating the 95% confidence interval. From the plot, we observe that: (1) in the single-charge scenario as shown in Figure 5(c) , Gear achieves a remarkably high $coverage$ score (over 0.95) with just 1 case retrieved. In the multi-charge scenario as shown in Figure 5(a) and Figure 5(b), Gear retrieve about 3 cases to encompass approximately 85% of the charges. These results demonstrate that Gear has considerable ability in charge prediction. This capability stems from our integration of judgment prediction into case retrieval, unifying the predictions for both tasks in a traversal on the law structure constraint tree. In other words, we not only require the model to retrieve relevant cases but also ensure that the retrieved cases share the same applicable charges as the query case. This incorporation brings the judgment prediction capability to Gear. (2) Gear demonstrates a significantly coverage improvement compared to DSI and DSI-QG on both datasets, especially when a limited number of cases are retrieved, such as 1 or 3. This mainly attributes to the specialized design of Gear for judgment prediction tasks including the law-aware hierarchical identifiers and the revision loss, which explicitly enhances the accuracy of judgment predictions. These experimental results verify Gear is capable of performing competitive legal judgment predictions.
3.5. Quality of Extracted Rationales
To answer RQ4, we conduct experiments and evaluate the rationales extracted by Gear in comparison to those generated by prevalent query generation methods featured in existing generative retrieval studies. We consider the Direct indexing (Tay et al., 2022), which means using the first 512 tokens as the query; Doc2query (Zhuang et al., 2022; Wang et al., 2022), which uses a language model to generate pseudo text as the model input; and ChatGPT, which demonstrates to provide text summaries that are on par with human-level quality (Zhang et al., 2023b; Bang et al., 2023; Liu et al., 2023b; Wang et al., 2023). First, we follow the practice (Zhao and Vydiswaran, 2021; Yu et al., 2022) and conduct the human evaluation to assess the quality of extracted rationales. We randomly selected 50 examples from both ELAM and LeCaRDv2 and ask two annotators with a legal background to determine whether the extracted rationales (generated queries) are sufficient to ascertain the applicable charges for the original cases, indicating that the extracted sentences comprehensively encompass the primary information from the original cases. Each annotator is provided with the rationales (generated queries) and the corresponding original case and is asked to mark the question as 1 if they agreed with it and 0 otherwise. As shown in Table 3, the results illustrate rationales extracted by Gear consistently outperform both Direct indexing and Doc2query on two datasets, achieving performance comparable to that of ChatGPT. These results confirm the rationales extracted by Gear is informative for both tasks.
Table 3. Time overhead and human evaluation of rationale extraction over 50 randomly sampled data from ELAM and LeCaRDv2 by two annotators with the inter-rater agreement of 0.96. “Time”(ms) denotes the extraction time per case.
ELAM LeCaRDv2
Methods Acc. Time Acc. Time
Direct indexing (Tay et al., 2022) 0.66 - 0.38 -
Doc2query (Zhuang et al., 2022) 0.48 4094.78(±1363.07) 0.26 3947.01(±1229.81)
ChatGPT 0.94 16332.84(±6201.80) 0.84 15750.02(±4853.38)
$f_{R}$ in Gear 0.94 0.10(±0.01) 0.86 0.67(±0.03)
Table 4. Comparison in average inference wall time (95% confidence interval) per query on a single Nvidia A6000.
Dataset
Models ELAM LeCaRDv2
Ultron 110.544(±24.703) 31.854(±7.766)
DSI-QG 73.147(±17.062) 26.369(±5.973)
DSI 72.646(±14.820) 27.023(±6.084)
Gear 59.184(±8.737) 20.391(±4.064)
Then, we also compare the time overhead of Gear for extracting rationales and baselines for generating queries. Please note that Direct indexing, as it directly takes the first 512 tokens as the query, has negligible time overhead. Besides, the runtime of ChatGPT API may be affected by network latency, we thus conduct 50 tests and report the average time overhead and its 95% confidence interval as shown in Table 3. From the results, we can see that Gear achieves performance similar to ChatGPT but with far less time consumption. Compared to Doc2query, Gear also exhibits a significant advantage in terms of time overhead.
To gain a more intuitive understanding of the differences among the texts generated by various methods, we conduct a case study in Appendix B.1.
4. Inference Time and Convergence
To answer RQ5, we record the inference wall time (in milliseconds) per query of baseline methods and Gear. This comparison is conducted on a single Nvidia A6000 across two datasets. From the results shown in Table 4, we observe that Gear demonstrates superior efficiency, with an inference time of 59.184 ms (±8.737) for the ELAM dataset, and 20.391 ms (±4.064) for LeCaRDv2. Specifically, when compared to Ultron, Gear achieves a remarkable 46.48% reduction in inference time for the ELAM dataset and a 36.03% reduction for LeCaRDv2. This is mainly because Ultron uses product quantization to create the identifier for each case. For all case embeddings, Ultron first divides embedding space into several groups and then performs $k$-means clustering on each group. It usually leads to excessively long identifiers. In the case of DSI and DSI-QG, Gear exhibits a substantial 19.01% and 18.59% improvement for ELAM and a 22.64% and 18.59% improvement for LeCaRDv2. The reason for this result is that in DSI and DSI-QG the tree constructed by $k$-means may be unbalanced, meaning that the lengths of case identifiers are unequal. Some case identifiers may be longer, which impairs the inference performance. These findings indicate the remarkable efficiency of Gear.
For further confirming Gear’s efficiency, we plot the testing curves for DSI, DSI-QG, and our Gear with the $x$-axis denoting the number of epochs, the $y$-axis denoting the MRR score, and shaded area indicating the 95% confidence interval. As shown in Figure 6, we observe that Gear not only outperforms DSI and DSI-QG significantly in terms of performance but also exhibits superior convergence. Gear achieves near-optimal MRR performance by 5 epochs, whereas DSI-QG and DSI converge at 6 and 8 epochs, respectively. The results verify the efficiency of Gear.
5. Related Work
5.1. Legal Case Retrieval
Legal case retrieval is a long-standing research topic in the field of information retrieval. In the early exploration, researchers (Moens, 2001; Zeng et al., 2005; Klein et al., 2006; Saravanan et al., 2009; Bench-Capon et al., 2012) made efforts to inject the legal knowledge to the case retrieval through the decomposition of legal issues and involving the ontology. In recent years, deep learning techniques have demonstrated their effectiveness in exploring the semantics of legal case retrieval. One representative line of works focused on the network-based precedents methods tailored for the common law systems. For example, Minocha et al. (2015) leveraged the Precedent Citation Network (PCNet) to predict the relevance based on whether the sets of precedent citations occur in the same cluster. Bhattacharya et al. (2020) proposed Hier-SPCNet to capture all domain information inherent in laws and precedents. The other line of works judged the relevance between the query case and candidate case according to their text-level similarity. One such method is BERT-PLI (Shao et al., 2020). It divided the legal case texts into several paragraphs and used BERT (Devlin et al., 2018) to obtain the similarity between the paragraphs, achieving promising ranking performance. Lawformer (Xiao et al., 2021b) is another text-based method. It used millions of Chinese criminal and civil case documents to pre-train a longformer (Beltagy et al., 2020) model and achieved impressive performance across legal tasks. To further estimate the relevance score in the legal domain, Li et al. (2023) made full use of the structure relationships of legal case documents for pre-training and achieved state-of-the-art performances.
Although these studies have made efforts to improve the retrieval performance, they ignored the charge-level consistency between the query case and the retrieved case, resulting in sub-optimal and unreliable results.
Figure 6. Testing curves of DSI, DSI-QG, and our Gear.
5.2. Generative Retrieval
Generative retrieval has recently emerged as a promising direction for document retrieval, enjoying merits like being end-to-end differentiable. Various methods have been introduced to generate semantic document identifiers. For example,  Cao et al. (2021) introduced a related sequence-to-sequence system to conduct entity retrieval. They first represented documents as unique names that are composed of entity names. Then they used a auto-regressive generation model to generate the unique names of these entities based on contextual information.  Tay et al. (2022) proposed the differentiable search index (DSI) paradigm, which is an auto-regressive generation model to perform ad-hoc retrieval tasks. The input of the model was a natural language query and the model regressively generated documents’ identifier strings that are relevant to the given query.  Wang et al. (2022) proposed a novel method NCI, which used a tailored prefix-aware weight-adaptive decoder to optimize the recall and ranking performance. NCI also used constrained beam search to make sure all the generated identifiers were valid. Ultron (Zhou et al., 2022) leveraged document titles and substrings as identifiers to enrich the semantic information of identifiers. To mitigate data distribution mismatch that occurs between the indexing and the retrieval phases,  Zhuang et al. (2022) proposed DSI-QG, which adopted a query generation model with a cross-encoder ranker to generate and select a set of relevant queries.
Existing studies focused on the general documents, lacking specific designs for legal cases, particularly the incorporation of law knowledge. As a result, it fails to deliver satisfactory performance.
6. Conclusion
In this study, we introduce Gear, a novel generative legal case retrieval method that integrates judgment prediction into the retrieval process. Gear explicitly exploits the law knowledge and extracts rationales from legal cases, ensuring a shared and informative representation for both tasks. Grounded with the inherent hierarchy of laws, Gear constructs a law structure constraint tree and assigns the law-aware semantic identifier to each case. These designs enable a unified traversal from the root, through intermediate charge nodes, to case-specific leaf nodes, which respectively correspond to thses two tasks. With the help of the proposed revision loss, Gear jointly minimizes the discrepancy between the identifiers of predicted and labeled charges/ retrieved cases, improving the accuracy and consistency for both tasks. Extensive experiments on two datasets demonstrate that Gear consistently outperforms state-of-the-art methods in legal case retrieval while maintaining competitive judgment prediction performance.
Appendix A REPRODUCIBILITY
A.1. Datasets
We conduct experiments on ELAM (Yu et al., 2022) and LeCaRDv2 (Ma et al., 2021) for legal case retrieval in this work.
ELAM (Yu et al., 2022) [^5] https://github.com/ruc-wjyu/OPT-Match. is a Chinese legal case retrieval dataset, focusing on criminal cases. ELAM has corresponding labels for both legal case retrieval and judgment prediction, which is suitable for our goal. We exclude those cases with multiple applicable charges to consider the retrieval and judgment prediction performance in a single charge scenario. The resulting candidate pool size of ELAM is 1332. Other data preprocessing is aligned to (Li et al., 2023).
LeCaRDv2666https://github.com/THUIR/LeCaRDv2. is the official updated version of LeCaRD  (Ma et al., 2021). In this dataset, the relevance labels are divided into four levels, ranging from 3 to 0, indicating a gradual decrease in relevance. We follow the data preprocessing approach of (Li et al., 2023), with the exception of increasing the candidate pool size from 100 to 1390 to further validate the effectiveness of baselines and our model. In LeCaRDv2, we consider both single and multiple charges, where each case corresponds 1.5 charges on average. Although LeCaRDv2 does not provide the judgment labels, we instead create pseudo labels using the proposed metric $coverage$ in Section 3.1.
A.2. Baselines
We consider three types of baselines in this study.
(1) Term-based methods:
  • Query Likelihood (QL) (Zhai et al., 2008) is a probabilistic language modeling approach used to estimate the relevance of documents to a given query. In QL, the likelihood of a document generating the words in a query is calculated, treating both the query and the document as random variables. This approach is based on the assumption that relevant documents are more likely to generate the query terms.
  • BM25 (Robertson et al., 2009) is a probabilistic information retrieval model widely used in the field of text retrieval. BM25 takes into account both term frequency and document length normalization, making it a robust and effective ranking model.
(2) Dense retrieval methods:
  • ANCE (Xiong et al., 2020) is a contrastive learning based model for ad-hoc dense retrieval. It selects hard training negatives from the entire corpus to reduce the stochastic gradient variance, and improve retrieval performance.
  • BERT+ANN (Devlin et al., 2018; Chalkidis et al., 2020) is a strong baseline in ad-hoc retrieval tasks that uses BERT to encode documents and then apply Approximate Nearest Neighbor (ANN) search algorithms to retrieve relevant document. In this work, we adopt the BERT pre-trained on a large Chinese legal corpus [^7] https://github.com/thunlp/OpenCLaP..
  • Lawformer+ANN (Xiao et al., 2021a): similar to BERT+ANN, we adopt Lawformer, which is a longformer (Beltagy et al., 2020) backbone and pre-trained on large legal case corpus, to encode legal cases.
  • SAILER (Li et al., 2023) is a structure-aware pre-training method for legal case retrieval. It adopts an asymmetric encoder-decoder architecture to integrate structures of legal case document information into dense vectors. SAILER is the state-of-the-art legal case retrieval method.
(3) Generative retrieval methods:
  • DSI (Tay et al., 2022) is a new paradigm for document retrieval tasks. It utilizes a Transformer-based encoder-decoder model to map queries directly to relevant identifiers. DSI achieves the end-to-end retrieval.
  • NCI (Wang et al., 2022) improves DSI in terms of using constrained beam search and prefix-aware weight-adaptive decoder.
  • DSI-QG (Zhuang et al., 2022) design a query generation process on the top of DSI, which can mitigate data distribution mismatches present between the indexing and the retrieval phases.
  • Ultron (Zhou et al., 2022) improves DSI through adopting product quantization to create semantic-based identifier and using URL to create term-based identifier.
A.3. Implementation Details
We implement baseline methods following the suggestions in the original papers. Specifically, (1) For classical term-based baselines, we use the pyserini [^8] https://github.com/castorini/pyserini. and genism [^9] https://github.com/RaRe-Technologies/gensim. toolkits with the default parameters. (2) For dense retrieval baselines, we use the open-sourced checkpoint [^10] https://github.com/thunlp. to initialize model parameters of BERT and Lawformer and use faiss [^11] https://github.com/facebookresearch/faiss.toolkit to implement ANN algorithms. The batch size is set to 16, the max length of input case is set to 512 for BERT and 1024 for Lawformer. We tune these models with in-batch contrastive loss. (3) For generative retrieval baselines, we directly use their official open-source implementations with the Chinese pre-trained T5 “Randeng”  [^12] https://huggingface.co/IDEA-CCNL/Randeng-T5-77M-MultiTask-Chinese. as backbone. We use beam search to retrieve relevant cases, where the beam size is set to 30.
We keep the backbone model and beam size same as baselines and set the max input length of Gear to 512, the rest hyperparameters are tuned as follows: the batch size is set to 2; the learning rate is tuned from [1e-5,1e-4] with step size 2e-5; in the rationale extraction module, $k_{1}$, $k_{2}$, $k_{3}$ are respectively set to 2, 5, 15 for ELAM and 10, 20, 15 for LeCaRDv2 ; $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ are respectively set to 10.0, 1.0, 0.1; in the law structure constraint tree module, the height of the tree (the length of the hierarchical identifier) $L$ is set to 4; in the training, $\mu$ is set to 1; $\gamma_{h}$ is tuned from $\{0.01,0.1\}$; $\gamma_{h}$ is set to 1; $\lambda_{w}$ is tuned from $\{1,10,100\}$; $\lambda_{l}$ is tuned from [1e-4,1e-2] with step size 5e-4. Hyperparameters of Gear are tuned using grid search with Adam(Kingma and Ba, 2014). ALL experiments are conducted on a single Nvidia A6000.
Figure 7. Comparison between the rationale extracted by Gear and the queries generated by Direct indexing, Doc2query, and ChatGPT. Underlined snippets denotes the key circumstances or key elements for the case.
Figure 8. A case study of prompting ChatGPT to obatin the key elements and key circumstances for a charge.
Appendix B Supplementary Experiments
B.1. Case Study in Rationale Extraction
In this section, we employ a case study to compare the performance of Gear and baseline methods in rationale extraction (query generation). Figure 7 showcases the comparison with all results translated from Chinese. We can find that (1) queries generated by Direct indexing and Doc2query fail to articulate the key circumstances and key elements of the original case, which are crucial factors for evaluating the relevance in legal case retrieval. This is because they lack law knowledge and struggle to handle the extremely lengthy documents. (2) The quality of both rationales extraction and ChatGPT summary is considerable as the underlined snippets highlighted. While both convey a substantial amount of key information, the rationales extracted by Gear are more sufficient and organized, which is easier for model learning. It contributes to our multi-level extraction strategy where we respectively extract keywords and sentences according to the definition of laws. The above results validate the effectiveness of the rationale extraction of Gear.
B.2. Building Corpus by Prompting ChatGPT
To improve the effectiveness of rationale extraction, we prompt chatGPT and extend the corpus using chatGPT’s feedback. Our prompt is shown in Figure 8. We would like ChatGPT play the role of a legal expert targeting on the Criminal Law and ask ChatGPT output the key circumstances and key elements for each charge. From the case, we found that chatGPT can provide multiple perspectives on a charge, including perpetrator, victim, behavior, demands, etc. These annotations serve as supplementary clarifications to the Criminal Law definitions and help Gear accurately select the informative rationales from legal cases.

### v2

@@
-Incorporating Judgment Prediction into Legal Case Retrieval via Law-aware Generative Retrieval
+Explicitly Integrating Judgment Prediction with Legal Document Retrieval: A Law-Guided Generative Approach
 Abstract.
-Legal case retrieval and judgment prediction are crucial components in intelligent legal systems.
-In practice, determining whether two cases share the same charges through legal judgment prediction is essential for establishing their relevance in case retrieval.
-However, current studies on legal case retrieval merely focus on the semantic similarity between paired cases, ignoring their charge-level consistency.
-This separation leads to a lack of context and potential inaccuracies in the case retrieval that can undermine trust in the system’s decision-making process.
-Given the guidance role of laws to both tasks and inspired by the success of generative retrieval, in this work, we propose to incorporate judgment prediction into legal case retrieval, achieving a novel law-aware Generative legal case retrieval method called Gear.
-Specifically, Gear first extracts rationales (key circumstances and key elements) for legal cases according to the definition of charges in laws, ensuring a shared and informative representation for both tasks.
-Then in accordance with the inherent hierarchy of laws, we construct a law structure constraint tree and assign law-aware semantic identifier(s) to each case based on this tree.
-These designs enable a unified traversal from the root, through intermediate charge nodes, to case-specific leaf nodes, which respectively correspond to two tasks.
-Additionally, in the training, we also introduce a revision loss that jointly minimizes the discrepancy between the identifiers of predicted and labeled charges as well as retrieved cases, improving the accuracy and consistency for both tasks.
-Extensive experiments on two datasets demonstrate that Gear consistently outperforms state-of-the-art methods in legal case retrieval while maintaining competitive judgment prediction performance.
-Legal case retrieval, Generative Retrieval, Legal Judgment Prediction
-copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
+Legal document retrieval and judgment prediction are crucial tasks in intelligent legal systems.
+In practice, determining whether two documents share the same judgments is essential for establishing their relevance in legal retrieval.
+However, existing legal retrieval studies either ignore the vital role of judgment prediction or rely on implicit training objectives, expecting a proper alignment of legal documents in vector space based on their judgments.
+Neither approach provides explicit evidence of judgment consistency for relevance modeling, leading to inaccuracies and a lack of transparency in retrieval.
+To address this issue, we propose a law-guided method, namely GEAR, within the generative retrieval framework.
+GEAR explicitly integrates judgment prediction with legal document retrieval in a sequence-to-sequence manner.
+Specifically, given the intricate nature of legal documents, we first extract rationales from documents based on the definition of charges in law.
+We then employ these rationales as queries, ensuring efficiency and producing a shared, informative document representation for both tasks.
+Second, in accordance with the inherent hierarchy of law, we construct a law structure constraint tree and represent each candidate document as a hierarchical semantic ID based on this tree.
+This empowers GEAR to perform dual predictions for judgment and relevant documents in a single inference, i.e., traversing the tree from the root through intermediate judgment nodes, to document-specific leaf nodes.
+Third, we devise the revision loss that jointly minimizes the discrepancy between the IDs of predicted and labeled judgments, as well as retrieved documents, thus improving accuracy and consistency for both tasks.
+Extensive experiments on two Chinese legal case retrieval datasets show the superiority of GEAR over state-of-the-art methods while maintaining competitive judgment prediction performance.
+Moreover, we validate the effectiveness of GEAR on a French statutory article retrieval dataset, reaffirming its robustness across languages and domains.
+Legal Document Retrieval, Generative Retrieval, Legal Judgment Prediction
+journalyear: 2024copyright: acmlicensedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657717isbn: 979-8-4007-0431-4/24/07ccs: Information systems Information retrievalccs: Applied computing Law
 1. Introduction
-Legal case retrieval, which seeks to retrieve relevant precedent cases based on a usually undecided case, stands as a pivotal component in intelligent legal systems.
-This task relies on specialized legal knowledge encoded in laws, with the goal of providing essential evidence to legal practitioners and supporting their decision-making processes.
-Existing studies on legal case retrieval (Ma et al., 2021; Xiao et al., 2021a; Chalkidis et al., 2020) have been focused on the semantic similarity between paired cases, for example,  Li et al.
-(2023) utilize the structural information contained in legal cases to conduct fine-grained semantic interactions between the query and candidate cases.
-Although they have achieved impressive performances, they ignore the fact that legal case retrieval and judgment prediction [^1] Legal judgment prediction aims to predict the relevant law article, charge, and term of penalty for a given query case.
-In this work, we focus on the charge prediction.
-Other sub-tasks are easily incorporated into our proposed method.
-are closely intertwined (Shao et al., 2023; Zhang and Dou, 2023).
-In particular, the charge-level consistency between the query and retrieved case, a direct product of the judgment prediction, is crucial for relevance estimation in legal case retrieval.
-As shown in Figure 1, it provides valuable contextual information and effective relevance signals.
-Moreover, incorporating judgment prediction into case retrieval improves the transparency (Olteanu et al., 2021; Kim et al., 2020; Baan et al., 2019) of the decision-making process.
-This integration aids legal practitioners in tracing the retrieved cases from the perspective of charges, thereby making the legal decisions trustworthy.
-Figure 1. A pair of relevant cases from the LeCaRDv2 dataset. The applicable charges (underlined) and their corresponding rationales (red and blue) provide effective relevance signals.
-Despite the aforementioned benefits, there remain the following challenges in incorporating judgment prediction into case retrieval.
-Firstly, legal case retrieval and judgment prediction are usually formulated as two distinct machine learning problems–ranking and classification.
-It is difficult for one ranking model to predict the applicable charges for legal cases and leverage the charge prediction to enhance ranking.
-Secondly, legal cases are inherently lengthy and complicated.
-As shown in Figure 1, only the rationales (key circumstances and key elements)  (Ma et al., 2021; Yu et al., 2022) of the case provide effective signals for relevance estimation.
-It is expected to represent each case as a shared and informative representation for both tasks, but existing methods often encode legal cases in a coarse-grained granularity way (Chalkidis et al., 2020; Xiao et al., 2021a).
-Thirdly, both tasks rely on specialized law knowledge (Sun et al., 2022; Liu et al., 2023c), an appropriate way that effectively incorporates law expertise to guide the prediction of both tasks remains a concern, given that current researches (Chalkidis et al., 2020; Xiao et al., 2021a; Zhang et al., 2023a; Li et al., 2023) solely focus on text-level semantics for predictions.
-Facing the above challenges and inspired by the success of generative retrieval (Tay et al., 2022; Wang et al., 2022), we propose a novel generative legal case retrieval method, namely Gear, to incorporate judgment prediction into legal case retrieval.
-As illustrated in Figure 2, Gear unifies these two tasks into a tree-based generative retrieval framework.
-The insight of Gear is to map the query case to a set of law-aware semantic identifiers where each identifier represents a precedent case relevant to the query and also reflects applicable charges.
-Specifically, we first construct a corpus based on the applicable laws [^2] We consider the Criminal Law of the People’s Republic of China in this work.
-and subsequently extract rationales at word- and sentence-level for each case according to this corpus.
-It effectively filters out the noise and ensures the case representations are shared and informative to both tasks.
-Then, we create the law structure constraint tree based on the inherent hierarchy of laws (Chapter-Section-Article) , considering that both tasks are learned with the guidance of laws.
-Thus, we can consider situations where each case may involve single or multiple charges and assign the law-aware semantic identifier(s) to each case given this tree, making the identifier reflect its charges.
-In this way, the retrieval is equivalent to a traversal from the root, through intermediate charge nodes, to case-specific leaf nodes, which respectively correspond to two tasks.
-To improve the accuracy and consistency for both tasks, we develop a revision loss that jointly minimizes the discrepancy between predicted and labeled charges/ retrieved cases.
-Furthermore, to assess the performance of judgment prediction in case retrieval, we introduce a novel metric called coverage.
-This metric gauges the charge-level consistency between the query and retrieved cases.
+Legal document retrieval and judgment prediction are fundamental components in intelligent legal systems.
+The former entails the retrieval of relevant legal documents (cases or statutory articles) give a query.
+On the other hand, the latter seeks to predict the outcomes or judgments rendered in legal cases, such as applicable charges, term-of-penalties, etc.
 Figure 1.
@@
 The middle part in blue is the generative retrieval framework.
+These two tasks are closely intertwined (Shao et al., 2023a, b; Zhang and Dou, 2023) in practice.
+From the retrieval side, determining whether two documents share the same judgments is essential for establishing their relevance.
+Regrettably, most of the existing studies (Ma et al., 2021; Xiao et al., 2021a; Chalkidis et al., 2020; Sun et al., 2024; Yu et al., 2022a) of legal document retrieval frequently overlook the significance of judgment prediction and merely focus on the text-level semantic similarity.
+Recently, Li et al.
+(2023a) introduced an implicit training objective that uses the fact description of the legal document to predict its judgment, expecting a proper alignment of legal documents in vector space based on their judgments.
+While these studies show effectiveness in retrieval performance, they fail to provide explicit evidence of judgment consistency for relevance modeling.
+Consequently, this limitation leads to inaccuracies and a lack of transparency (Olteanu et al., 2021; Kim et al., 2020; Baan et al., 2019).
+It is because their legal relevance reasoning especially regarding judgment remains unclear, and we cannot trace back the decision-making process based on the retrieval results.
+Therefore, we aim to explicitly integrate judgment prediction with legal document retrieval.
+However, there remain the following challenges to achieve our goal.
+Firstly, legal document retrieval and judgment prediction are usually formulated as two distinct machine learning problems–retrieval and classification.
+It is difficult for one retrieval model to predict the applicable judgment for legal cases and in turn leverage the judgment prediction to enhance retrieval.
+Secondly, legal documents are inherently lengthy and complicated.
+It results in the retrieval efficiency issue and hiders to represent each document as a shared and informative representation for both tasks.
+Thirdly, both tasks rely on specialized law knowledge (Sun et al., 2022; Liu et al., 2023b), an appropriate way that effectively injects law expertise to guide the prediction of both tasks remains a concern.
+Facing the above challenges, we propose a novel law-guided generative legal document retrieval method, namely GEAR.
+GEAR explicitly integrates judgment prediction with legal document retrieval in a sequence-to-sequence (Seq2Seq) manner as illustrated in Figure 1.
+The insight of GEAR lies in formulating the retrieval process into the generation of law-aware semantic IDs, where each ID not only represents a document relevant to the query but also reflects its applicable judgments.
+Specifically, we first construct a corpus based on the definition of charges in law [^1] In this paper, we take laws in China and Belgium as examples.
+It is worth noting that GEAR can readily accommodate laws from various other countries.
+and subsequently extract rationales (Yu et al., 2022b; Sun et al., 2023) representing the key elements/ circumstances (Ma et al., 2021; Shao et al., 2023b) at the word- and sentence-level for each document according to this corpus.
+We employ these rationales instead of raw documents as queries for generation.
+This strategy not only serves to effectively filter out the noise of legal documents, ensuring generation efficiency but also renders the rationales shared and informative for both tasks.
+Then, we create the law structure constraint tree based on the inherent hierarchy of law (e.g.
+Chapter-Section-Article), considering that both tasks are learned with the guidance of law.
+Given this tree, we assign legal documents hierarchical semantic IDs, with the IDs reflecting their judgments, for example, the ID “0-2-5-269-809” indicates the document named 809 falls under Article 269 of Chapter 5 of Section 2.
+In this way, the generation of these IDs is equivalent to traversing the tree from the root through intermediate judgment nodes, to document-specific leaf nodes.
+It makes GEAR capable of showing the legal reasoning process and performing dual predictions for judgment and relevant documents in a single inference.
+To further improve the accuracy and consistency of both tasks, we devise a novel training objective called the revision loss.
+This loss aligns with the hierarchy of the tree and jointly minimizes the discrepancy between predicted and labeled judgments/ retrieved cases.
+Extensive experiments on two Chinese legal case retrieval datasets show the consistent superiority of GEAR over state-of-the-art methods while maintaining competitive judgment prediction performance.
+We also validate the effectiveness of GEAR on a French statutory article retrieval dataset, reaffirming its generalization ability.
 The major contributions of the paper are summarized as follows:
-(1) To the best of our knowledge, this is the first work that incorporates judgment prediction into legal case retrieval, addressing the challenge of charge-level consistency between the query case and retrieved cases in legal search.
-Our work also improves the transparency of the legal decision-making process.
-(2) We propose a novel model, namely Gear, to unify legal case retrieval and judgment prediction under the generative retrieval framework.
-We explicitly leverage the law knowledge to extract rationales from cases, assign cases the law-aware hierarchical identifies, and formulate the prediction as a traversal on the law structure constraint tree.
-We also propose a revision loss to jointly improve the accuracy and consistency of both tasks.
-(3) We conduct extensive experiments on two public datasets and propose an evaluation metric for assessing the charge-level consistency between the query and retrieved cases.
-Experimental results indicate Gear not only achieves state-of-the-art performance in legal case retrieval but also maintains competitive judgment prediction performance.
+(1) To the best of our knowledge, this is the first work that explicitly integrates judgment prediction with legal document retrieval.
+Our method is capable of showing the legal reasoning process and performing dual predictions for both tasks in a single inference.
+It improves the transparency of the legal decision-making.
+(2) We propose a novel law-guided generative model, namely GEAR.
+We explicitly leverage the law knowledge to extract rationales from legal documents, assign them the law-aware hierarchical IDs, and formulate the prediction as a traversal on the law structure constraint tree.
+We also propose the revision loss to jointly improve the accuracy and consistency of both tasks.
+(3) We conduct extensive experiments on three public datasets of legal document retrieval in two languages.
+The results indicate GEAR not only achieves state-of-the-art performance in legal case and statutory article retrieval but also maintains competitive judgment prediction performance.
+2. Related Work
+2.1. Legal Document Retrieval
+Legal document retrieval is a long-standing research topic in the field of information retrieval.
+In the early exploration, researchers (Moens, 2001; Zeng et al., 2005; Klein et al., 2006; Saravanan et al., 2009; Bench-Capon et al., 2012) made efforts to inject the legal knowledge to the retrieval through the decomposition of legal issues and involving the ontology.
+In recent years, deep learning demonstrated its effectiveness in exploring document semantics.
+One representative line of works focused on the network-based precedents methods tailored for the common law systems.
+For example, Minocha et al.
+(2015) leveraged the Precedent Citation Network (PCNet) to predict the relevance based on whether the sets of precedent citations occur in the same cluster.
+Bhattacharya et al.
+(2020) proposed Hier-SPCNet to capture all domain information inherent in laws and precedents.
+The other line of works judged the relevance between the query and candidate document according to their text-level similarity.
+One such method was BERT-PLI (Shao et al., 2020).
+It divided the legal document into several paragraphs and used BERT (Devlin et al., 2018) to obtain the similarity between the paragraphs.
+Lawformer (Xiao et al., 2021b) was another text-based method.
+It used millions of Chinese criminal and civil case documents to pre-train a Longformer (Beltagy et al., 2020) model.
+Despite impressive performance, these works overlook the significance of judgment prediction and merely focus on the text-level similarity, resulting in sub-optimal and unreliable results.
+Recently, Li et al.
+(2023a) introduced an implicit training objective that uses the fact description of the legal document to predict its judgment, expecting a proper alignment of legal documents in vector space based on their judgments.
+It fails to provide explicit evidence of judgment consistency for relevance modeling, leading to inaccuracies and a lack of transparency.
+2.2. Generative Retrieval
+Generative retrieval has recently emerged as a promising direction for document retrieval.
+These methods assign semantic IDs to documents and utilize language models for ID generation.
+It enables an end-to-end retrieval in contrast to the traditional index-then-retrieve paradigm.
+Various methods have been introduced to generate semantic document IDs.
+For example,  Cao et al.
+(2021) introduced a Seq2Seq system to conduct entity retrieval.
+They first represented documents as unique names that are composed of entity names.
+Then they used a auto-regressive generation model to generate the unique names of these entities based on contextual information.
+Tay et al.
+(2022) proposed the differentiable search index (DSI) paradigm, which is an auto-regressive generation model to perform ad-hoc retrieval tasks.
+The input of the model was a natural language query and the model regressively generated documents’ ID strings that are relevant to the given query.
+Wang et al.
+(2022) proposed a novel method NCI, which used a tailored prefix-aware weight-adaptive decoder to optimize the retrieval performance.
+Ultron (Zhou et al., 2022) leveraged document titles and substrings as IDs to enrich the semantic information of IDs.
+To mitigate data distribution mismatch that occurs between the indexing and the retrieval phases,  Zhuang et al.
+(2022) proposed DSI-QG, which adopted a query generation model with a cross-encoder to generate and select a set of relevant queries.
+Existing studies focused on the general domain, lacking specific designs for legal documents and the integration of law knowledge.
 3. Methodology
 3.1. Task Formulation
-In this work, we target on the task of legal case retrieval.
+In this work, we target on legal document retrieval including legal case retrieval (LCR) and statutory article retrieval (SAR).
 Suppose that we have a set of collected samples $\mathcal{D}=\{(q,\mathcal{C},\mathcal{R})\}$.
-For each data instance, $q$ represents a undecided query case submitted by the legal practitioner; $\mathcal{C}=\{c_{1},c_{2},\cdots,c_{N}\}$ represents candidate cases (precedents) pool with size $N\in\mathbb{N}^{+}$; $\mathcal{R}$ represents the labeled relevant case set from $\mathcal{C}$ given the query case $q$.
-Unlike previous studies that only predict to retrieve $\mathcal{R}$ from $\mathcal{C}$ given $q$, in this work, we instead unify judgment prediction and case retrieval into a generative retrieval framework, and thus aims at learning a ranking function $f:q\times\mathcal{C}\rightarrow\mathcal{R}\times\mathcal{E}$, where $\mathcal{E}$ denotes the set of applicable charges corresponding to $q$.
-2.2. The Principle of Our Method
-To retrieve relevant cases and predict their applicable charges, we develop a novel learning method from the viewpoint of generative retrieval (GR).
-Essentially, GR is an emerging new retrieval paradigm, that adopts a deep neural network to perform the end-to-end retrieval in contrast to traditional the index-then-retrieve paradigm.
-GR models consist of two major steps to directly generate identifier strings of documents as the retrieval target.
-In the first indexing step that focuses on memorizing the information about each document, GR models take each original document $c$ as input and generate its identifier ${id}^{c}$ as output in a straightforward Seq2Seq fashion.
+For each data instance, $q$ is the query representing an undecided legal case submitted by the legal practitioner in LCR, a legal question in SAR; $\mathcal{C}=\{c_{1},c_{2},\cdots,c_{N}\}$ with size $N\in\mathbb{N}^{+}$ is the candidate precedent case set in LCR, the statutory article pool in SAR; $\mathcal{R}$ represents the labeled relevant case/ statutory article set from $\mathcal{C}$ given the query $q$.
+Unlike previous studies that only predict to retrieve $\mathcal{R}$ from $\mathcal{C}$ given $q$, in this work, we instead unify judgment prediction and document retrieval into a generative retrieval framework, and thus aim at learning a retrieval function $f:q\times\mathcal{C}\rightarrow\mathcal{R}\times\mathcal{E}$, where $\mathcal{E}$ denotes the set of applicable judgment corresponding to $q$.
+3.2. Overall Framework
+To learn $f$ and explicitly integrate judgment prediction with legal document retrieval, we develop GEAR, a novel law-guided generative approach from the viewpoint of generative retrieval.
+Essentially, given a query document, GEAR adopts a language model to perform the Seq2Seq generation where the retrieved documents are represented as semantic IDs.
+Following the practice of (Tay et al., 2022; Wang et al., 2022; Zhuang et al., 2022; Zhou et al., 2022), GEAR consists of two major steps to directly generate IDs of documents as the retrieval target.
+In the first indexing step that focuses on memorizing the information about each document, our GEAR takes each document $c$ as input and generates its ID ${id}^{c}$ as output.
 The model is trained with the standard language model objective with the teacher forcing:
-(1) $\mathcal{L}_{i}=\sum_{c\in\mathcal{C}}\log P({id}^{c}|c),$
-where ${id}^{c}$ is represented by the atomic identifier or the semantic identifier.
-Atomic identifiers are generated by assigning each document an arbitrary integer, while semantic identifiers are generated according to the hierarchical clustering algorithm employed over all the documents.
-In the second retrieval phase, GR models associate each $q$ to its relevant document $r$ through an auto-regressive generation:
+(1) $\mathcal{L}_{i}=\sum_{c\in\mathcal{C}}\log P({id}^{c}|c).$
+In the second retrieval phase, GEAR models associate each $q$ to its relevant document $r$ through an auto-regressive generation:
 (2) $\mathcal{L}_{r}=\sum_{(q,r)\in\mathcal{D}}\log P({id}^{r}|q),$
-where ${id}^{r}$ denotes the identifier of $r$.
-${id}^{r}$ is also represented by the atomic identifier or the semantic identifier.
-As such, once a GR model is learned, it can be used to retrieve candidate documents for a test query in an end-to-end manner using beam search.
-As aforementioned in Section 1, applying GR to our legal scenario faces several challenges including the lengthy and complicated content, the law-aware semantic identifier, and the training objective for enhancing both tasks.
-Therefore, when incorporating legal judgment prediction into case retrieval in the above GR framework, the critical learning tasks become: (1) extract rationales instead of using original documents to form the input for GR (Section 2.3); (2) create informative law-aware identifiers for each case based on the hierarchical structure of laws (Section 2.4); (3) develop a training objective to explicitly ensure charge-level and case-level consistency between predictions and labels(Section 2.5).
+where ${id}^{r}$ denotes the ID of $r$. As such, once a GEAR model is trained, it can be used to retrieve candidate documents for a test query in an end-to-end manner using beam search.
+As aforementioned in Section 1, there are several challenges in the legal domain.
+For explicitly integrating judgment prediction with legal document retrieval within GEAR, the critical learning tasks become: (1) extract rationales instead of using raw documents to form the input for generation (Section 3.3); (2) create informative law-aware IDs for each document based on the hierarchical structure of law (Section 3.4); (3) develop a training objective to explicitly ensure judgment-level and document-level consistency between predictions and labels (Section 3.5).
 Figure 2.
 The proposed three modules of GEAR.
-$f_{R}$ extract rationales from legal cases; $f_{T}$ assigns hierarchical identifiers to each case and constrain the decoding; $\mathcal{L}_{c}$ jointly optimizes case retrieval and judgement prediction.
+$f_{R}$ extract rationales from legal documents; $f_{T}$ assigns hierarchical IDs to each document and constrain the decoding; $\mathcal{L}_{c}$ jointly optimizes document retrieval and judgement prediction.
 3.3. Rationale Extraction
-To filter the noise and provide shared and informative representations for legal cases, based on the set of laws $L$, we devise a module called $f_{R}$ to extract rationales $E$ instead of using the original document $doc\in\{q,c\}$ as the input of Gear:
+To ensure efficiency and provide shared and informative representations for legal documents, based on the set of laws $L$, we devise a module called $f_{R}$ to extract rationales $E$ instead of using the raw document $doc\in\{q,c\}$ as the input of GEAR:
 (3) $E=f_{R}(doc,L),$
 where $E=\{E_{w},E_{s}\}$ in which $E_{w}$ and $E_{s}$ respectively denotes the word-level and sentence-level rationales.
-As shown in Figure 3, to achieve this goal, we first leverage the guidance of laws to construct a corpus $B$, collecting law-based keywords and embeddings.
-Specifically, the corpus consists of three parts $B=\{B_{1},B_{2},B_{3}\}$.
-$B_{1}$ is constructed according to the lexical variants of all charges names in $L$, for example, in terms of the charge “crime of forges the seals of a company, enterprise, institution or a people’s organization”(translated from Chinese), we split the charge name and put all keywords and their lexical variants to the corpus after the stop words are removed.
-$B_{2}$ is constructed similarly to the $B_{1}$.
-We put all keywords in the definition of laws and the feedback from prompting ChatGPT [^3] https://openai.com/chatgpt, the prompts are provided in Appendix B.2.
-to the corpus after the stop words are removed.
-$B_{3}$ is constructed by collecting BERT (Devlin et al., 2018; Chalkidis et al., 2020) embeddings [^4] https://github.com/thunlp/OpenCLaP.
-for the definition of each law.
-Once the corpus is collected, we compute the multiple-level scores and extract $E$ for each case based on three corpora as follows.
-At the word level, we split the case and respectively select the top-$k_{1}$ (top-$k_{2}$) keywords $E_{w_{1}}$ ($E_{w_{2}}$) from $B_{1}$ ($B_{2}$) as follows:
+As illustrated in Figure 2, we first leverage the guidance of law to construct a corpus $B=\{B_{1},B_{2},B_{3}\}$, collecting law-based keywords $B_{1},B_{2}$ and embeddings $B_{3}$.
+The keyword set $B_{1}$ is constructed according to the lexical variants of all charge names in $L$.
+For example, in terms of the charge “crime of forges the seals of a company, enterprise, institution or a people’s organization”(translated from Chinese), we split the charge name and remove the stop words, then add the rest and their lexical variants to $B_{1}$.
+The keyword set $B_{2}$ is constructed similarly to the $B_{1}$.
+We split the definitions of charges in law and remove the stop words then add the rest and their lexical variants to $B_{2}$.
+For the augmentation purpose, we employ the definitions of charges in law as prompts for the large language model [^2] We use ChatLaw (Cui et al., 2023) for Chinese data.
+Since we have not found a suitable LLM for the Belgian legal domain, we omit this step for this data.
+(LLM) designed for the legal domain.
+After obtaining feedback from LLM, we remove stop words and incorporate the results into $B_{2}$.
+To avoid the hallucination issue, we engage legal experts to manually assess the quality of augmentations to ensure the effectiveness of $B_{2}$.
+$B_{3}$ is constructed by collecting Legal-BERT (Devlin et al., 2018; Chalkidis et al., 2020) embeddings for the definition of each charge in law.
+Once the corpus is collected, we compute the multiple-level scores and extract $E$ for each document based on three corpora as follows.
+At the word level, we split the document and respectively select the top-$k_{1}$ (top-$k_{2}$) keywords $E_{w_{1}}$ ($E_{w_{2}}$) from $B_{1}$ ($B_{2}$) as follows:
 (4) $E_{w_{i}}=\mathop{\arg\text{top-}k_{i}}_{w\in B_{i}}\biggl{(}\text{tf}(w,doc)\biggr{)},$
@@
 where $w$ is the word from $s$.
-Please note that since query cases are typically undecided with no charges, we extract rationales using the method described above.
-On the other hand, candidate cases are precedents with decided charges, thus, when extracting the rationales for them, we shrunk $B$ into a corpus constructed based on its corresponding charges.
+Please note that since query documents are typically undecided i.e., without labeled applicable charges, we extract rationales using the method described above.
+For the candidate precedent documents, whose applicable charges are given, we shrunk $B$ into a corpus constructed based on their corresponding labeled charges.
 3.4. Law Structure Constraint Tree
-Considering that both legal case retrieval and judgment prediction require guidance by the law where there usually exists a “Chapter-Section-Article” hierarchy, we argue that the decision process to judge whether two cases are relevant in case retrieval is analogous to the search in such tree-like hierarchy.
-In other words, when the practitioners search for the relevant cases given a query case, they always expect the charge of the relevant precedent case to be located at the same position in the hierarchy as the charges applicable in the query case.
-Therefore, we devise a module $f_{T}$ that leverages the inherent hierarchy of the law to construct a law structure constraint tree $T$ as shown in Figure 3 and assigns the law-aware semantic identifier $id$ for each $c\in\mathcal{C}$:
+Given that both legal document retrieval and judgment prediction require guidance by law, typically organized in a “Chapter-Section-Article” hierarchy, we argue that the decision process to judge whether two documents are relevant in document retrieval is analogous to the search in such a tree-like hierarchy.
+In other words, when legal practitioners search for the relevant documents given a query, they always expect the charge of relevant documents to be located at the same position in the law hierarchy as the charges applicable in the query.
+Therefore, we devise a module $f_{T}$ that leverages the inherent hierarchy of the law to construct a law structure constraint tree $T$ as illustrated in Figure 2 and assigns the law-aware semantic ID $id$ for each $c\in\mathcal{C}$:
 (7) $id=f_{T}(c,T),$
@@
 The prefix depends on the position of the applicable charges within the tree.
-As illustrated in Figure 3, case 809 involves the charge of “crime of robbery” which falls under Article 269 of Chapter 5 of Section 2 of the Criminal Law of the People’s Republic of China.
-Hence, the assigned prefix for this case is 0-2-5-269, where 0 represents the root node of the tree.
-Regarding the suffix, we regard legal cases are the children of their corresponding charges on the tree, and assign a unique identifier to each case under a crime node.
-Compared to current works that employ hierarchical $k$-means to create identifiers for each document, ours avoid that same integers may have different meanings at different levels, and thus ensure the effectiveness of model learning.
-In the retrieval (inference), the identifiers of top relevant cases can be easily generated via beam search.
-Due to the hierarchical nature of laws as analyzed above, it is convincing to use constrained beam search (Anderson et al., 2017; Hu et al., 2019; Post and Vilar, 2018) decoding process with our law structure constraint tree, which in turn only generates the valid identifiers.
-On the other hand, one unique feature of the legal domain is that a single case can involve multiple charges.
-When the query is this kind of case, the ideal retrieval results should encompass these charges.
-Therefore, we assign $k$ identifiers to cases involving $k$ charges, with each identifier corresponding to a specific charge.
-For example, as shown in Figure 3, the case 725 involves charge 267 and charge 269, the valid identifiers for this case include 0-2-5-267-725 and 0-2-5-269-725.
-In doing so, during the retrieval, we expect the model to retrieve all identifiers of the target case, thereby increasing the probability of the target being retrieved.
+As shown in Figure 2, document 809 involves the charge of “crime of robbery” which falls under Article 269 of Chapter 5 of Section 2 of the Criminal Law of the People’s Republic of China.
+Hence, the assigned prefix for this document is 0-2-5-269, where 0 represents the root node of the tree.
+As for the suffix, we regard documents are the children of their corresponding charges on the tree, and assign a unique ID to each of them under a crime node.
+Compared to current works that employ hierarchical $k$-means to create IDs for each document, ours avoid same integers may have different meanings at different levels, and thus ensure the effectiveness of model training.
+On the other hand, one unique feature of the legal domain is that a single document can involve multiple charges.
+When the query is this kind of document, the ideal retrieval results should encompass these charges.
+Therefore, we assign $k$ IDs to documents involving $k$ charges, with each ID corresponding to a specific charge.
+For example, as shown in Figure 2, the 725 involves charge 267 and charge 269, the valid IDs for this document include 0-2-5-267-725 and 0-2-5-269-725.
+In doing so, during retrieval, we expect the model to retrieve all IDs of the target documents, thereby increasing the probability of the target being retrieved.
 3.5. Revision Loss
-Besides employing the typical language model training objectives (Equation 1 and Equation 2), we also develop a revision loss for the consistency between the query and retrieved cases, i.e., we aim at directly minimizing their charge-level and case-level discrepancy.
-Formally, as illustrated in Figure 3, given the predicted identifier (list of integers) $[\hat{id}_{1},\hat{id}_{2},\cdots,\hat{id}_{L}]$ and the corresponding ground-truth identifier $[{id}_{1},{id}_{2},\cdots,{id}_{L}]$ for a query $q$ at the same position of the ranking, both having a length of $L$, we compare the difference between them and calculate the reward $R_{t}$ at each step $t\in[1,L]$ as follows:
+Besides employing the typical language model training objectives (Equation 1 and Equation 2), we also develop a novel training objective called the revision loss for the consistency between the query and retrieved documents, i.e., we aim to directly minimize their judgment-level and document-level discrepancy.
+Formally, as illustrated in Figure 2, given the predicted ID (list of integers) $[\hat{id}_{1},\hat{id}_{2},\cdots,\hat{id}_{L}]$ and the corresponding ground-truth ID $[{id}_{1},{id}_{2},\cdots,{id}_{L}]$ for a query $q$, both having a length of $L$, we compare the difference between them and calculate the reward $R_{t}$ at each step $t\in[1,L]$ as follows:
 (8) $R_{t}=\begin{cases}\mu,&\text{if}~{}\hat{id}t={id}_{t},\\ -\gamma_{h}^{L-t}\mu|\hat{id}_{t}-{id}_{t}|,&\text{if}~{}\hat{id}_{t}\neq{id}_{t},\\ \end{cases}$
-where $\mu$ is the constant reward unit, $\gamma_{h}\in(0,1]$ is the hierarchy penalty factor used to penalize the differences between predictions and labels layer by layer along the law structure constraint tree, with larger penalties as it get closer to the tree root and smaller penalties as it get closer to the tree leaves.
-Intuitively, if two cases share the same prefix, they are likely relevant to each other because of the same applicable charge, thus receiving higher rewards.
-Conversely, if two cases belong to different chapters in terms of applicable charges, they are likely dissimilar resulting in lower rewards.
+where $\mu$ is the constant reward unit, $\gamma_{h}\in(0,1]$ is the hierarchy penalty factor used to penalize the differences between predictions and labels layer by layer along the law structure constraint tree, with larger penalties as it gets closer to the tree root and smaller penalties as it gets closer to the tree leaves.
+Intuitively, if two documents share the same prefix, they are likely relevant to each other because of the same applicable charge, receiving higher rewards.
 Then we apply the REINFORCE algorithm (Williams, 1992) to optimize the model parameters, the revision loss is defined to minimize the policy gradient objective:
 (9) $\mathcal{L}_{c}=-\sum_{q\in\mathcal{D}}\sum^{bz}_{b}\gamma_{r}^{b}\sum^{L}_{t}\log p(\hat{id}_{t}|q)\cdot R_{t}-\lambda_{w}\log p({id}_{t}|q),$
-where $\gamma_{r}\in(0,1]$ the optional rank penalty factor used to focus on the top of ranking; $bz$ denotes the beam size, i.e., the number of cases to retrieve for each query; $p(\hat{id}_{t}|q)$ is the probability that predicting to generate $\hat{id}_{t}$ given $q$ at layer $t$; $\lambda_{w}$ is the balance coefficient.
-To further handling the sparse reward issue and improve the training efficiency, we follow (Xin et al., 2020; Chen et al., 2021; Liu et al., 2023a) and add the second term to Equation 9 that directly increases the probability of generating ${id}_{t}$.
-Therefore, the overall training objective for Gear is:
+where $\gamma_{r}\in(0,1]$ the optional penalty factor used to focus on the top of retrieved document list; $bz$ denotes the beam size, i.e., the number of documents to retrieve for each query; $p(\hat{id}_{t}|q)$ is the probability that predicting to generate $\hat{id}_{t}$ given $q$ at layer $t$; $\lambda_{w}$ is the balance coefficient.
+To further handle the sparse reward issue and improve the training efficiency, we follow (Xin et al., 2020; Chen et al., 2021; Liu et al., 2023a) and add the second term to Equation 9 that directly increases the probability of generating ${id}_{t}$.
+Thus, the overall training objective is:
 (10) $\mathcal{L}=\mathcal{L}_{i}+\mathcal{L}_{r}+\lambda_{l}\mathcal{L}_{c},$
 where $\lambda_{l}$ is the coefficient to balance the indexing loss (Equation 1), retrieval loss (Equation 2), and the revision loss.
-2.6. Discussion
-In this part, we highlight the merits and limitations of the proposed Gear approach as follows:
-Transparency.
-Inspired by the decision-making process in legal scenarios, our approach formulates the legal case retrieval as the traversal on the law structure constraint tree.
-This design not only provides valuable contextual information but also improves the transparency of retrieval results by tracing back the decision-making process.
-Versatility.
-Since the retrieved cases are represented by the law-aware semantic identifier with each digit denoting one prediction at each layer in the law structure constraint tree, we can directly predict the applicable judgments for the query case according to the retrieval results.
-This design makes it versatile to unify legal case retrieval and judgment prediction.
-Limitation.
-In this work, we mainly focus on the Criminal Law of the People’s Republic of China.
-It follows the civil law system and has an explicit hierarchy.
-In countries that follow the common law system such as the United States and the United Kingdom, their laws usually do not possess an explicit hierarchical structure.
-It may limit the applicability of our Gear.
-Fortunately, for example, in the United States, there is a combination of federal and state laws, where the latter often follows a structure involving “titles, chapters, sub-chapters” hierarchy.
-Therefore, our method can be applied to such scenarios.
+3.6. Inference
+In the inference, we aim to retrieve the top-$k$ documents from the candidate pool.
+Since we have assigned hierarchical semantic ID to each document based on the law-aware constraint tree where each leaf node corresponds to a candidate document in the pool, we utilize the constrained beam search (Anderson et al., 2017; Hu et al., 2019; Post and Vilar, 2018) to ensure all the generated document IDs are valid within the tree.
 4. Experiments
-In this section, we conduct experiments to answer the following research questions:
-  • RQ1: How does Gear perform on legal case retrieval compared to state-of-the-art methods?
+In this section, we conduct experiments to answer the following research questions: RQ1: How does GEAR perform on legal document retrieval compared to state-of-the-art methods?
 RQ2: How effective are the three modules in GEAR?
-  • RQ3: Can Gear show competitive performance on judgment prediction?
+RQ3: Can GEAR show competitive performance on judgment (applicable charges) prediction?
 RQ4: What is the quality of the rationales extracted by GEAR including effectiveness and efficiency?
-  • RQ5: Can Gear incur less time overhead in legal case retrieval compared to popular generative methods?
+RQ5: Can GEAR incur less time overhead in legal document retrieval compared to popular generative methods?
+RQ6: How robust is GEAR across languages and domains (e.g.
+in statutory article retrieval)?
 4.1. Experimental Settings
 4.1.1. Datasets
-We conduct experiments on LeCaRDv2(Ma et al., 2021) and ELAM(Yu et al., 2022) for legal case retrieval.
-Table 1 reports basic statistics of both datasets.
-We leave the detailed descriptions in Appendix A.1.
-Table 1. Statistics of ELAM and LeCaRDv2. “(Top 1)” indicates the charges corresponding to the most relevant candidate case serve as the label charges for the query case.
+ELAM [^3] https://github.com/ruc-wjyu/OPT-Match.
+(Yu et al., 2022b) is a Chinese LCR dataset, focusing on criminal cases.
+ELAM has corresponding labels for both case retrieval and judgment prediction, which is suitable for our goal.
+We exclude those cases with multiple applicable charges to consider the retrieval and judgment prediction performance in a single charge scenario.
+The resulting candidate pool size of ELAM is 1332.
+Other data preprocessing is aligned to (Li et al., 2023a).
+LeCaRDv2444https://github.com/THUIR/LeCaRDv2.
+(Li et al., 2023b) is the official updated version of LeCaRD  (Ma et al., 2021).
+In this dataset, the relevance labels are divided into four levels, ranging from 3 to 0, indicating a gradual decrease in relevance.
+We follow the data preprocessing approach of (Li et al., 2023a), with the exception of increasing the candidate pool size from 100 to 1390 to further validate the effectiveness of baselines and our model.
+In LeCaRDv2, cases encompass both single and multiple charges, averaging 1.5 charges per case.
+Considering the ground-truth judgment labels of query cases have not been provided, we ask two legal experts (Ph.D.
+in Law) to annotate the charge label for the testing queries.
+The experts are proficient in Chinese criminal law with sufficient experience in handling cases similar to this dataset.
+They carefully align the LeCaRDv2 judgment criteria (Li et al., 2023b) before annotation and discuss opinions to reach a consensus, ensuring accurate labeling.
+BSARD [^5] https://huggingface.co/datasets/maastrichtlawtech/bsard.
+(Louis and Spanakis, 2021) is a SAR dataset composed of more than 1.1K legal questions labeled by domain experts with relevant articles selected from the 22K law articles gathered from 32 publicly available Belgian codes.
+It is worth noting that BSARD contains structural annotations of corresponding laws, facilitating the utilization of law structural knowledge.
+4.1.2. Baselines
+We consider three types of baselines in this study.
+(1) Sparse retrieval methods: Query Likelihood (QL) (Zhai et al., 2008) is a probabilistic language modeling approach employed to assess the relevance of documents to a provided query.
+BM25 (Robertson et al., 2009) is a probabilistic information retrieval model widely used in the field of text retrieval.
+BM25 takes into account both term frequency and document length normalization.
+(2) Dense retrieval methods: BERT (Devlin et al., 2018) is a strong baseline in ad-hoc retrieval tasks in the open domain.
+In this paper, we adopt the checkpoint that is pre-trained on a large Chinese corpus Following (Tay et al., 2022; Wang et al., 2022; Zhuang et al., 2022), after encoding legal documents using BERT, we then apply Approximate Nearest Neighbor (ANN) search algorithms to retrieve relevant documents.
+Legal-BERT [^6] https://github.com/thunlp/OpenCLaP.
+(Chalkidis et al., 2020) is a variant of BERT that undergoes specific training in the legal domain to better understand and process text related to law.
+Lawformer  (Xiao et al., 2021a) is a Longformer (Beltagy et al., 2020) backbone pre-trained on large legal case corpus, to encode legal texts.
+ChatLaw-Text2Vec [^7] https://huggingface.co/chestnutlzj/ChatLaw-Text2Vec.
+(Cui et al., 2023) is a legal text matching model based on ChatLaw which is pre-trained on a corpus of 936,727 legal documents.
+G-DSR (Louis et al., 2023) uses legal-CamemBERT [^8] https://huggingface.co/maastrichtlawtech/legal-camembert-base., a legal variant of CamemBERT trained on BSARD dataset.
+It takes into account both the dense representation of text and the graph representation of legal structures.
+G-DSR is the state-of-the-art SAR method in the French legal domain.
+SAILER  (Li et al., 2023a) is a structure-aware LCR model.
+It adopts an asymmetric encoder-decoder architecture to integrate structures of legal case document information into dense vectors.
+SAILER achieves state-of-the-art retrieval performance in Chinese LCR domain.
+As for the training of all baselines, we follow (Tay et al., 2022; Wang et al., 2022; Zhuang et al., 2022) and use Approximate nearest neighbor Negative Contrastive Estimation (ANCE) (Xiong et al., 2020) method.
+(3) Generative retrieval methods: DSI (Tay et al., 2022) is a new paradigm for document retrieval tasks.
+It utilizes a Transformer-based encoder-decoder model to map queries directly to relevant IDs.
+DSI achieves the end-to-end retrieval.
+NCI (Wang et al., 2022) improves DSI in terms of using constrained beam search and prefix-aware weight-adaptive decoder.
+Both DSI and NCI use hierarchical $k$-means clustering of document vectors to create $k$-means IDs.
+DSI-QG (Zhuang et al., 2022) design a query generation process on the top of DSI, which can mitigate data distribution mismatches present between the indexing and the retrieval phases.
+Ultron (Zhou et al., 2022) improves DSI through adopting product quantization to create semantic IDs and using URLs to create term-based IDs.
+4.1.3. Implementation Details
+We implement baseline methods following the suggestions in the original papers.
+(1) For classical term-based baselines, we use the pyserini and genism toolkits with the default parameters.
+(2) For dense retrieval baselines, we use the open-sourced checkpoint to initialize model parameters of the pre-trained models and use faiss toolkit to implement ANN algorithms.
+The batch size is set to 16.
+The max length of the input text is set to 1024 for Lawformer and 512 for the other models.
+We tune these models with in-batch contrastive loss.
+(3) For generative retrieval baselines, we directly use their official open-source implementations, employing the pre-trained T5 “Randeng” [^9] https://huggingface.co/IDEA-CCNL/Randeng-T5-77M-MultiTask-Chinese.
+as backbone for Chinese legal domain, and “t5-base” [^10] https://huggingface.co/t5-base.
+Please note that this is applicable to French datasets.
+for the French legal domain.
+We use beam search to retrieve relevant cases, where the beam size is set to 30.
+We keep the backbone model and beam size same as baselines and set the max input length of GEAR to 512, the rest hyperparameters are tuned as follows: the batch size is set to 2; the learning rate is tuned from [1e-5,1e-4] with step size 2e-5; in the rationale extraction module, $k_{1}$, $k_{2}$, $k_{3}$ are respectively set to 2, 5, 15 for ELAM and 10, 20, 15 for LeCaRDv2; $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ are respectively set to 10.0, 1.0, 0.1; in the law structure constraint tree module, the height of the tree (the length of the hierarchical ID) $L$ is set to 4; in the training, $\mu$ is set to 1; $\gamma_{h}$ is tuned from $\{0.01,0.1\}$; $\gamma_{h}$ is set to 1; $\lambda_{w}$ is tuned from $\{1,10,100\}$; $\lambda_{l}$ is tuned from [1e-4,1e-2] with step size 5e-4.
+Hyperparameters of GEAR are tuned using grid search with Adam(Kingma and Ba, 2014).
+ALL experiments are conducted on a single NVIDIA RTX A6000.
+The source code and datasets have been shared at: https://github.com/E-qin/GEAR.
+Table 1. Statistics of the datasets. Since queries in BSARD do not involve judgments, we omit the judgment prediction comparison on this dataset.
 Dataset
-Statistics ELAM LeCaRDv2
-Avg. length per candidate case 1163.68 1568.38
-Avg. length per query case 1304.98 558.18
-Avg. # charges per candidate case 1.00 1.50
-Avg. # charges per query case (Top 1) 1.00 1.53
-# Train candidate cases/query 1332 1390
-# Query cases involved 147 653
-# Crimes involved 97 100
-[…]
+Statistics ELAM LeCaRDv2 BSARD
+Avg. length per candidate document 1163.68 1568.38 880.29
+Avg. length per query document 1304.98 558.18 92.48
+Avg. # charges per candidate case 1.00 1.50 -
+# Available candidates per query 1332 1390 1612
+# Query documents involved 147 653 1108
+# Charges involved of judgment 97 100 -
+Table 2.
+Performance comparisons of our approach and the baselines on ELAM dataset and LeCaRDv2 dataset.
+The best and the second-best performances are denoted in bold and underlined fonts, respectively.
+“R@K” is short for “Recall@K”.
+denotes GEAR performs signifcantly better than baselines based on two-tailed paired t-test with Bonferroni correction ($p<0.05$).
 ELAM LeCaRDv2
@@
 BM25 (Robertson et al., 2009) 0.0340 0.0680 0.1497 0.2245 0.0635 0.0435 0.1452 0.2549 0.3900 0.1862
-ANCE (Xiong et al., 2020) 0.0769 0.1894 0.2357 0.3300 0.1512 0.0362 0.0929 0.0968 0.1145 0.1323
-BERT (Devlin et al., 2018)+ANN 0.0302 0.0908 0.1405 0.2461 0.0921 0.0299 0.0718 0.0757 0.0934 0.1162
-Lawformer (Xiao et al., 2021a)+ANN 0.0484 0.1638 0.2209 0.3185 0.1410 0.0318 0.0591 0.0728 0.1093 0.1238
-SAILER (Li et al., 2023) 0.0670 0.2132 0.3282 0.4604 0.1729 0.0208 0.0844 0.1510 0.2371 0.1318
+BERT (Devlin et al., 2018) 0.0302 0.1008 0.1405 0.2861 0.1521 0.0299 0.0718 0.1557 0.2534 0.1162
+Legal-BERT (Zhong et al., 2019) 0.0384 0.0938 0.1509 0.3123 0.1510 0.0218 0.0620 0.1081 0.2743 0.1138
+Lawformer (Xiao et al., 2021a) 0.0537 0.1682 0.2220 0.3789 0.1701 0.0518 0.1491 0.2728 0.3593 0.1638
+ChatLaw-Text2Vec (Cui et al., 2023) 0.0385 0.1371 0.2065 0.3323 0.1694 0.0356 0.0813 0.1510 0.3380 0.1379
+SAILER (Li et al., 2023a) 0.0729 0.2132 0.3282 0.4604 0.2029 0.0608 0.1644 0.2910 0.4271 0.2018
 DSI (Tay et al., 2022) 0.0204 0.1134 0.2274 0.3159 0.1249 0.0232 0.0577 0.0768 0.1285 0.1159
@@
 Ultron (Zhou et al., 2022) 0.0607 0.1583 0.2260 0.3506 0.1678 0.0333 0.1207 0.2142 0.3492 0.1511
-Gear ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
+GEAR 0.0793 0.2368 0.3356 0.4976 0.2365 0.0630 0.1706 0.3142 0.4625 0.2162
    w/o $\mathcal{L}_{c}$ 0.0611 0.1657 0.2793 0.4167 0.1763 0.0452 0.1485 0.2549 0.4478 0.1979
@@
    w/o $f_{T}$ 0.0464 0.1429 0.2328 0.3158 0.1626 0.0166 0.0550 0.0971 0.1388 0.1596
-3.1.2. Baselines
-We compare Gear with three types of representative baselines in this work including term-based methods, dense retrieval methods, and generative retrieval methods.
-We leave the introduction on baselines and the detailed implementation in Appendix A.2 and Appendix A.3, respectively.
 4.1.4. Evaluation Metrics
-Following previous works (Tay et al., 2022; Zhuang et al., 2022; Wang et al., 2022; Zhou et al., 2022), we leverage the commonly adopted metrics, including Recall (R) and MRR.
+For a fair comparison, we follow previous works (Tay et al., 2022; Zhuang et al., 2022; Wang et al., 2022; Zhou et al., 2022) and leverage the commonly adopted metrics, including Recall (R) and MRR.
 The averaged results on all test cases are reported.
-To demonstrate Gear’s ability on judgment prediction, we introduce a new metric called $coverage@k$, assessing the percentage of charges involved in the query cases that are covered by top-$k$ retrieved cases.
-This metric evaluates the charge-level consistency between the query case and the retrieved cases.
-Considering that query cases are usually undecided and LeCaRDv2 does not have the ground-truth judged charges for query cases, we adopt $coverage_{\text{freq}}$ and $coverage_{\text{top1}}$ to create a pseudo label for each query case.
-In terms of $coverage_{\text{freq}}$, for each query case, we create a charge dictionary from all the charges involved in all relevant candidate cases and select the charge that appears most frequently as the charge label for the query.
-In terms of $coverage_{\text{top1}}$, for each query case, we take the charge involved in the most relevant candidate case as the charge label for the query.
+To demonstrate GEAR’s ability on judgment prediction, we introduce a new metric called $coverage@k$, assessing the percentage of charges involved in the query that are covered by top-$k$ retrieved documents.
+This metric evaluates the charge-level consistency, i.e., the extent to which the retrieval process contributes to the efficacy of legal judgment.
+Formally, $coverage@k$ is defined as:
+(11) $coverage@k=\frac{1}{N_{q}}\sum_{i=1}^{N_{q}}\frac{|\mathcal{E}_{i}^{\text{Top-}k}\cap\mathcal{E}_{i}|}{|\mathcal{E}_{i}|},$
+where |||~{}|| | is the size of a set; for $i$-th testing query, $\mathcal{E}_{i}$ denotes its label set of applicable judgment charges, $\mathcal{E}_{i}^{\text{Top-}k}$ denotes the the set of charges of the top-$k$ retrieved documents; $N_{q}$ is the number of testing queries.
 In our experiments, we compute $coverage@k$ metric at 1, 3, 5, and 10.
@@
 From the results, we have the following observations for RQ1:
-$\bullet$ Gear demonstrates a significant performance advantage over all baseline methods on both datasets.
-The relative improvements in terms of MRR on the ELAM and LeCaRDv2 are at least 36.77% and 16.09%, respectively.
+(1) GEAR demonstrates a significant performance advantage over all baseline methods on both datasets.
+The relative improvements of MRR on the ELAM and LeCaRDv2 are at least 16.55% and 7.13%, respectively.
 These results indicate GEAR’s effectiveness.
-We contribute the improvement to the elaborate design for incorporating judgment into case retrieval including rationale extraction, identifier construction, and the revision loss.
-See Section 3.3 for detailed analysis for each module of Gear.
-$\bullet$ Compared to sparse retrieval and dense retrieval methods, current generative retrieval methods struggle to achieve satisfying performance in the legal domain.
-Generative retrieval baselines merely focus on mapping original cases to their $k$-means style identifiers.
-Without injecting the law knowledge, these methods inevitably yield sub-optimal results.
-SAILER uses the case-level structure and obtains the second-best performance in ELAM where both query and candidate cases have clear structure.
-In LeCaRDv2 where only the fact part is provided for query case and there exhibits a distribution shift between query and candidate, BM25 performs second best.
-In contrast, Gear explicitly leverages the law knowledge to extract rationales, construct identifiers, and involve judgment prediction for legal case retrieval.
-These designs lead to significant performance improvement.
-3.3. Ablation Studies
+We attribute the improvement to the elaborate design for explicitly integrating judgment into case retrieval including rationale extraction, law-aware ID assignment, and the revision loss.
+See Section 4.3 for the detailed analysis of each module of GEAR.
+(2) Compared to sparse retrieval and dense retrieval methods, current generative retrieval methods struggle to achieve satisfying performance in the legal domain.
+Without injecting the law knowledge, generative retrieval baselines inevitably yield sub-optimal results.
+On the other hand, SAILER introduces an implicit training objective that uses the fact description of the legal document to predict its judgments.
+In this way, SAILER produces judgment-aware document representations and achieves the second-best performance.
+However, SAILER fails to provide explicit evidence of judgment consistency for relevance modeling, leading to an inferior retrieval performance compared to GEAR.
 Figure 3.
@@
 The means values of 5 repeated experiments are reported, with error bars representing the 95% confidence interval of the means.
-To answer RQ2, we conduct an ablation study to investigate the impact of each component of Gear.
-Firstly, we test the performance of Gear’s variants by removing a certain component.
-Following DSI (Tay et al., 2022), we replace rationale extraction and law-aware hierarchical identifiers to the Direct indexing (the first 512 tokens) and $k$-means identifiers.
-As shown in the bottom part of Table 2, we observe that: (1) Removing modules $\mathcal{L}_{c}$, $f_{R}$, and $f_{T}$ individually results in a performance degradation of 25.51%, 23.83%, and 31.16% on ELAM, 8.50%, 25.00%, and 26.13% on LeCaRDv2.
-These results verify the effectiveness of all three modules.
-(2) It is worth noting that removing our law-aware hierarchical identifiers results in the most significant performance decrease.
-It is because we inject law knowledge into each digit of the identifier, aligning them with the hierarchical structure of laws.
-By simulating the retrieval to mirror the legal decision-making process, we enable Gear to effectively learn the association between query cases and candidate cases.
-These results demonstrate the effectiveness of our identifiers and highlight the importance of introducing the law knowledge to legal case retrieval.
-To further validate the efficacy of the components of Gear, we combine the three components of Gear with DSI individually and test the retrieval performance.
-We conduct 5 repeated experiments and present the mean values of MRR, with error bars representing the 95% confidence interval.
-As illustrated in Figure 4, incorporating our components of Gear into DSI consistently results in performance improvements.
-Specifically, DSI equipped with rationale extraction achieves an improvement of over 60% on ELAM, and DSI equipped with law-aware identifiers demonstrated a 42% improvement on LeCaRDv2 dataset.
-However, DSI equipped with revision loss does not exhibit a substantial improvement.
-Because DSI utilizes $k$-means identifiers and the length of these identifiers may vary due to differences in cluster sizes.
-In such cases, our revision loss cannot accurately measure the differences between predictions and labels, resulting in a slight performance improvement.
-(a) $coverage_{\text{freq}}$ on LeCaRDv2
-(b) $coverage_{\text{top1}}$ on LeCaRDv2
-(c) $coverage$ on ELAM
+(a) $coverage$ on ELAM
+(b) $coverage$ on LeCaRDv2
 Figure 4.
@@
 The proposed GEAR consistently outperforms DSI and DSI-QG in both single (ELAM) and multiple (LeCaRDv2) charges scenarios.
+4.3. Ablation Studies
+To answer RQ2, we conduct an ablation study to investigate the impact of each component of GEAR.
+Firstly, we test the performance of GEAR’s variants by removing a certain component.
+Following DSI (Tay et al., 2022), we replace rationale extraction and law-aware hierarchical IDs with Direct indexing (using the first 512 tokens of document as queries for language models) and $k$-means IDs.
+As shown in the bottom part of Table 2, we observe that: (1) Removing modules $\mathcal{L}_{c}$, $f_{R}$, and $f_{T}$ individually results in a performance degradation of 25.51%, 23.83%, and 31.16% on ELAM, 8.50%, 25.00%, and 26.13% on LeCaRDv2.
+These results verify the effectiveness of all three modules.
+(2) It is worth noting that removing our law-aware hierarchical IDs results in the most significant performance decrease.
+It is because we inject law knowledge into each digit of the ID, aligning them with the hierarchical structure of laws.
+By simulating the retrieval to mirror the legal decision-making process, we enable GEAR to effectively learn the association between query cases and candidate cases.
+These results demonstrate the effectiveness of our IDs and highlight the importance of introducing structural and semantic knowledge in law to legal document retrieval.
+To further validate the efficacy of the components of GEAR, we combine the three components of GEAR with DSI individually and test its retrieval performance.
+We conduct 5 repeated experiments and present the mean values of MRR, with error bars representing the 95% confidence interval.
+As illustrated in Figure 3, integrating the components of GEAR into DSI consistently results in performance improvements.
+Specifically, DSI equipped with rationale extraction achieves an improvement of over 60% on ELAM, and DSI equipped with law-aware IDs demonstrated a 42% improvement on LeCaRDv2 dataset.
+However, DSI equipped with revision loss does not exhibit a substantial improvement.
+Because DSI utilizes $k$-means IDs and the length of these IDs may vary due to differences in cluster sizes.
+In such cases, our revision loss cannot accurately measure the differences between predictions and labels, resulting in a slight performance improvement.
 4.4. Quality of Judgment Prediction
-To answer RQ3, we conduct experiments to evaluate the judgment prediction accuracy of retrieved cases using the proposed metric $coverage$.
-Considering that in LeCaRDv2, the ground-truth judgment label of query cases is not provided, we use $coverage_{\text{freq}}$ and $coverage_{\text{top1}}$ to create pseudo labels for them.
+To answer RQ3, we conduct experiments to evaluate the judgment prediction accuracy of retrieved cases using the proposed metric $coverage@k$.
+Please note that we consider the applicable charges as the judgment results.
 We run experiments ten times and present the results in Figure 4, reporting the average score with the shaded area indicating the 95% confidence interval.
-From the plot, we observe that: (1) in the single-charge scenario as shown in Figure 5(c) , Gear achieves a remarkably high $coverage$ score (over 0.95) with just 1 case retrieved.
-In the multi-charge scenario as shown in Figure 5(a) and Figure 5(b), Gear retrieve about 3 cases to encompass approximately 85% of the charges.
+From the plot, we observe that: (1) in the single-charge scenario (ELAM) as shown in Figure 4(a) , GEAR achieves a remarkably high $coverage$ score (over 0.95) with just 1 case retrieved.
+In the multi-charge scenario (LeCaRDv2) as shown in Figure 4(b), GEAR retrieves about 3 cases to encompass approximately 85% of the charges.
 These results demonstrate that GEAR has considerable ability in charge prediction.
 This capability stems from our integration of judgment prediction into case retrieval, unifying the predictions for both tasks in a traversal on the law structure constraint tree.
-In other words, we not only require the model to retrieve relevant cases but also ensure that the retrieved cases share the same applicable charges as the query case.
-This incorporation brings the judgment prediction capability to Gear.
 (2) GEAR demonstrates a significantly coverage improvement compared to DSI and DSI-QG on both datasets, especially when a limited number of cases are retrieved, such as 1 or 3.
-This mainly attributes to the specialized design of Gear for judgment prediction tasks including the law-aware hierarchical identifiers and the revision loss, which explicitly enhances the accuracy of judgment predictions.
-These experimental results verify Gear is capable of performing competitive legal judgment predictions.
-3.5. Quality of Extracted Rationales
+This mainly attributes to the specialized design of GEAR for judgment prediction tasks including the law-aware hierarchical IDs and the revision loss, which explicitly enhances the accuracy of judgment predictions.
+These results verify GEAR is capable of performing competitive legal judgment predictions.
+4.5. Effectiveness and Efficiency of Rationale Extraction
 To answer RQ4, we conduct experiments and evaluate the rationales extracted by GEAR in comparison to those generated by prevalent query generation methods featured in existing generative retrieval studies.
-We consider the Direct indexing (Tay et al., 2022), which means using the first 512 tokens as the query; Doc2query (Zhuang et al., 2022; Wang et al., 2022), which uses a language model to generate pseudo text as the model input; and ChatGPT, which demonstrates to provide text summaries that are on par with human-level quality (Zhang et al., 2023b; Bang et al., 2023; Liu et al., 2023b; Wang et al., 2023).
-First, we follow the practice (Zhao and Vydiswaran, 2021; Yu et al., 2022) and conduct the human evaluation to assess the quality of extracted rationales.
-We randomly selected 50 examples from both ELAM and LeCaRDv2 and ask two annotators with a legal background to determine whether the extracted rationales (generated queries) are sufficient to ascertain the applicable charges for the original cases, indicating that the extracted sentences comprehensively encompass the primary information from the original cases.
-Each annotator is provided with the rationales (generated queries) and the corresponding original case and is asked to mark the question as 1 if they agreed with it and 0 otherwise.
-As shown in Table 3, the results illustrate rationales extracted by Gear consistently outperform both Direct indexing and Doc2query on two datasets, achieving performance comparable to that of ChatGPT.
-These results confirm the rationales extracted by Gear is informative for both tasks.
+We consider the following baselines: Direct indexing (Tay et al., 2022), which means using the first 512 tokens of the raw document as the query; Doc2query (Zhuang et al., 2022; Wang et al., 2022), which uses a language model to generate pseudo text as the model input; and ChatLaw (Cui et al., 2023), which demonstrates to provide legal summaries that are on par with human-level quality.
+First, we follow the practice (Zhao and Vydiswaran, 2021; Yu et al., 2022b) and conduct the human evaluation to assess the quality of extracted rationales.
+We randomly selected 50 samples from both ELAM and LeCaRDv2 and asked two annotators (Ph.D.
+in Law) to determine whether the extracted rationales (generated queries) are sufficient to ascertain the applicable charges for the original cases, indicating that the extracted sentences comprehensively encompass the primary information from the original cases.
+Each annotator is provided with a pair of the rationales (generated queries) and the corresponding original case for each sample and is asked to mark the sample as 1 if they agreed with it and 0 otherwise.
+We calculate the accuracy of the sample rationales based on annotators’ evaluation.
+As shown in Table 3, the results illustrate rationales extracted by GEAR consistently outperform both Direct indexing and Doc2query on two datasets, achieving performance comparable to that of ChatLaw.
+These results confirm the rationales extracted by GEAR are informative for both tasks.
+Then, we conduct an experiment to validate the time overhead of rationale extraction.
+Since Direct indexing takes the first 512 tokens as the query, we omit its time overhead.
+From the results shown in Table 4, we can see that GEAR achieves impressive accuracy on par with ChatLaw but with far less time consumption.
+Compared to Doc2query, GEAR also exhibits a significant advantage in terms of time overhead.
+These results verify the GEAR’s efficiency in terms of rationale extraction.
+Based on the results, the generative approaches for rationales are not advisable for legal document retrieval due to the efficiency issue.
 Table 3.
-Time overhead and human evaluation of rationale extraction over 50 randomly sampled data from ELAM and LeCaRDv2 by two annotators with the inter-rater agreement of 0.96.
-“Time”(ms) denotes the extraction time per case.
+Time overhead and human evaluation of rationale extraction over 50 random samples from ELAM and LeCaRDv2 by two annotators with the inter-rater agreement of 0.96.
+“Time”(ms) denotes the extraction time per sample.
 ELAM LeCaRDv2
@@
 Doc2query (Zhuang et al., 2022) 0.48 4094.78(±1363.07) 0.26 3947.01(±1229.81)
-ChatGPT 0.94 16332.84(±6201.80) 0.84 15750.02(±4853.38)
+ChatLaw (Cui et al., 2023) 0.94 20975.92(+3858.29) 0.88 21827.05(±5820.58)
 $f_{R}$ in GEAR 0.94 0.10(±0.01) 0.86 0.67(±0.03)
-Table 4. Comparison in average inference wall time (95% confidence interval) per query on a single Nvidia A6000.
+Table 4. Comparison in average inference wall time (95% confidence interval) per query on a single NVIDIA RTX A6000.
 Dataset
@@
 GEAR 59.184(±8.737) 20.391(±4.064)
-Then, we also compare the time overhead of Gear for extracting rationales and baselines for generating queries.
-Please note that Direct indexing, as it directly takes the first 512 tokens as the query, has negligible time overhead.
-Besides, the runtime of ChatGPT API may be affected by network latency, we thus conduct 50 tests and report the average time overhead and its 95% confidence interval as shown in Table 3.
-From the results, we can see that Gear achieves performance similar to ChatGPT but with far less time consumption.
-Compared to Doc2query, Gear also exhibits a significant advantage in terms of time overhead.
-To gain a more intuitive understanding of the differences among the texts generated by various methods, we conduct a case study in Appendix B.1.
-4. Inference Time and Convergence
+Table 5.
+Performance comparisons of our approach and the baselines on BSARD.
+“R@K” is short for “Recall@K”.
+indicates that improvements are significant based on two-tailed paired t-test with Bonferroni correction ($p<0.05$).
+Models R@5 R@10 R@20 R@30 MRR
+QL (Zhai et al., 2008) 0.1752 0.2128 0.2698 0.2923 0.1770
+BM25 (Robertson et al., 2009) 0.1815 0.2381 0.2871 0.3056 0.1786
+G-DSR (Louis et al., 2023) 0.1636 0.3426 0.5081 0.6720 0.3012
+DSI (Tay et al., 2022) 0.4578 0.5536 0.6315 0.6562 0.3955
+GEAR 0.5297 0.6164 0.7031 0.7170 0.4081
+4.6. Inference Time and Convergence
 To answer RQ5, we record the inference wall time (in milliseconds) per query of baseline methods and GEAR.
-This comparison is conducted on a single Nvidia A6000 across two datasets.
+This comparison is conducted on a single NVIDIA RTX A6000 across two datasets.
 From the results shown in Table 4, we observe that GEAR demonstrates superior efficiency, with an inference time of 59.184 ms (±8.737) for the ELAM dataset, and 20.391 ms (±4.064) for LeCaRDv2.
 Specifically, when compared to Ultron, GEAR achieves a remarkable 46.48% reduction in inference time for the ELAM dataset and a 36.03% reduction for LeCaRDv2.
-This is mainly because Ultron uses product quantization to create the identifier for each case.
-For all case embeddings, Ultron first divides embedding space into several groups and then performs $k$-means clustering on each group.
-It usually leads to excessively long identifiers.
+This is mainly because Ultron uses product quantization to create the IDs for documents.
+For all document embeddings, Ultron first divides embedding space into several groups and then performs $k$-means clustering on each group.
+It usually leads to excessively long IDs.
 In the case of DSI and DSI-QG, GEAR exhibits a substantial 19.01% and 18.59% improvement for ELAM and a 22.64% and 18.59% improvement for LeCaRDv2.
-The reason for this result is that in DSI and DSI-QG the tree constructed by $k$-means may be unbalanced, meaning that the lengths of case identifiers are unequal.
-Some case identifiers may be longer, which impairs the inference performance.
-These findings indicate the remarkable efficiency of Gear.
-For further confirming Gear’s efficiency, we plot the testing curves for DSI, DSI-QG, and our Gear with the $x$-axis denoting the number of epochs, the $y$-axis denoting the MRR score, and shaded area indicating the 95% confidence interval.
-As shown in Figure 6, we observe that Gear not only outperforms DSI and DSI-QG significantly in terms of performance but also exhibits superior convergence.
-Gear achieves near-optimal MRR performance by 5 epochs, whereas DSI-QG and DSI converge at 6 and 8 epochs, respectively.
+In DSI and DSI-QG, the tree constructed by $k$-means may be unbalanced, meaning that the lengths of case IDs are unequal.
+Some case IDs may be longer, which impairs the inference performance.
+For further confirming GEAR’s efficiency, we plot the testing curves for DSI, DSI-QG, and our GEAR with the $x$-axis denoting the number of epochs, the $y$-axis denoting the MRR score, and the shaded area indicating the 95% confidence interval.
+As illustrated in Figure 5, we observe that GEAR not only outperforms DSI and DSI-QG significantly in terms of performance but also exhibits superior convergence.
+GEAR achieves near-optimal MRR performance by 6 epochs, whereas DSI-QG and DSI converge at 8 and 10 epochs, respectively.
 The results verify the efficiency of GEAR.
-5. Related Work
-5.1. Legal Case Retrieval
-Legal case retrieval is a long-standing research topic in the field of information retrieval.
-In the early exploration, researchers (Moens, 2001; Zeng et al., 2005; Klein et al., 2006; Saravanan et al., 2009; Bench-Capon et al., 2012) made efforts to inject the legal knowledge to the case retrieval through the decomposition of legal issues and involving the ontology.
-In recent years, deep learning techniques have demonstrated their effectiveness in exploring the semantics of legal case retrieval.
-One representative line of works focused on the network-based precedents methods tailored for the common law systems.
-For example, Minocha et al.
-(2015) leveraged the Precedent Citation Network (PCNet) to predict the relevance based on whether the sets of precedent citations occur in the same cluster.
-Bhattacharya et al.
-(2020) proposed Hier-SPCNet to capture all domain information inherent in laws and precedents.
-The other line of works judged the relevance between the query case and candidate case according to their text-level similarity.
-One such method is BERT-PLI (Shao et al., 2020).
-It divided the legal case texts into several paragraphs and used BERT (Devlin et al., 2018) to obtain the similarity between the paragraphs, achieving promising ranking performance.
-Lawformer (Xiao et al., 2021b) is another text-based method.
-It used millions of Chinese criminal and civil case documents to pre-train a longformer (Beltagy et al., 2020) model and achieved impressive performance across legal tasks.
-To further estimate the relevance score in the legal domain, Li et al.
-(2023) made full use of the structure relationships of legal case documents for pre-training and achieved state-of-the-art performances.
-Although these studies have made efforts to improve the retrieval performance, they ignored the charge-level consistency between the query case and the retrieved case, resulting in sub-optimal and unreliable results.
 Figure 5. Testing curves of DSI, DSI-QG, and our GEAR.
-5.2. Generative Retrieval
-Generative retrieval has recently emerged as a promising direction for document retrieval, enjoying merits like being end-to-end differentiable.
-Various methods have been introduced to generate semantic document identifiers.
-For example,  Cao et al.
-(2021) introduced a related sequence-to-sequence system to conduct entity retrieval.
-They first represented documents as unique names that are composed of entity names.
-Then they used a auto-regressive generation model to generate the unique names of these entities based on contextual information.
-Tay et al.
-(2022) proposed the differentiable search index (DSI) paradigm, which is an auto-regressive generation model to perform ad-hoc retrieval tasks.
-The input of the model was a natural language query and the model regressively generated documents’ identifier strings that are relevant to the given query.
-Wang et al.
-(2022) proposed a novel method NCI, which used a tailored prefix-aware weight-adaptive decoder to optimize the recall and ranking performance.
-NCI also used constrained beam search to make sure all the generated identifiers were valid.
-Ultron (Zhou et al., 2022) leveraged document titles and substrings as identifiers to enrich the semantic information of identifiers.
-To mitigate data distribution mismatch that occurs between the indexing and the retrieval phases,  Zhuang et al.
-(2022) proposed DSI-QG, which adopted a query generation model with a cross-encoder ranker to generate and select a set of relevant queries.
-Existing studies focused on the general documents, lacking specific designs for legal cases, particularly the incorporation of law knowledge. As a result, it fails to deliver satisfactory performance.
+4.7. Robustness across Languages and Domains
+To answer RQ6, we compared the performance of baselines and GEAR on BSARD, a French SAR dataset.
+Since BSARD has a relatively short query length, we omit the rationale extraction part and take the raw legal questions as queries for GEAR.
+In terms of the ID, we follow the structure of the Belgian code provided in the data and assign the hierarchical semantic to each statutory article.
+From the results shown in Table 5, we have two observations: (1) with a small number of documents to retrieve, generative retrieval methods (DSI and GEAR) exhibit significantly higher retrieval performances in SAR scenario compared to dense retrieval.
+We assume the reason why dense models perform poorer is that there exists a significant gap between legal questions and statutory articles.
+It is difficult for dense models to learn the correct association between them especially without the law knowledge injected.
+(2) GEAR demonstrates the best performance, exhibiting a substantial advantage over sparse and dense retrieval methods including the current state-of-the-art model G-DSR.
+The improvement of GEAR benefits from explicitly injecting legal knowledge into generative retrieval frameworks.
 5. Conclusion
-In this study, we introduce Gear, a novel generative legal case retrieval method that integrates judgment prediction into the retrieval process.
-Gear explicitly exploits the law knowledge and extracts rationales from legal cases, ensuring a shared and informative representation for both tasks.
-Grounded with the inherent hierarchy of laws, Gear constructs a law structure constraint tree and assigns the law-aware semantic identifier to each case.
-These designs enable a unified traversal from the root, through intermediate charge nodes, to case-specific leaf nodes, which respectively correspond to thses two tasks.
-With the help of the proposed revision loss, Gear jointly minimizes the discrepancy between the identifiers of predicted and labeled charges/ retrieved cases, improving the accuracy and consistency for both tasks.
-Extensive experiments on two datasets demonstrate that Gear consistently outperforms state-of-the-art methods in legal case retrieval while maintaining competitive judgment prediction performance.
-Appendix A REPRODUCIBILITY
-A.1. Datasets
-We conduct experiments on ELAM (Yu et al., 2022) and LeCaRDv2 (Ma et al., 2021) for legal case retrieval in this work.
-ELAM (Yu et al., 2022) [^5] https://github.com/ruc-wjyu/OPT-Match.
-is a Chinese legal case retrieval dataset, focusing on criminal cases.
-ELAM has corresponding labels for both legal case retrieval and judgment prediction, which is suitable for our goal.
-We exclude those cases with multiple applicable charges to consider the retrieval and judgment prediction performance in a single charge scenario.
-The resulting candidate pool size of ELAM is 1332.
-Other data preprocessing is aligned to (Li et al., 2023).
-LeCaRDv2666https://github.com/THUIR/LeCaRDv2.
-is the official updated version of LeCaRD  (Ma et al., 2021).
-In this dataset, the relevance labels are divided into four levels, ranging from 3 to 0, indicating a gradual decrease in relevance.
-We follow the data preprocessing approach of (Li et al., 2023), with the exception of increasing the candidate pool size from 100 to 1390 to further validate the effectiveness of baselines and our model.
-In LeCaRDv2, we consider both single and multiple charges, where each case corresponds 1.5 charges on average.
-Although LeCaRDv2 does not provide the judgment labels, we instead create pseudo labels using the proposed metric $coverage$ in Section 3.1.
-A.2. Baselines
-We consider three types of baselines in this study.
-(1) Term-based methods:
-  • Query Likelihood (QL) (Zhai et al., 2008) is a probabilistic language modeling approach used to estimate the relevance of documents to a given query.
-In QL, the likelihood of a document generating the words in a query is calculated, treating both the query and the document as random variables.
-This approach is based on the assumption that relevant documents are more likely to generate the query terms.
-  • BM25 (Robertson et al., 2009) is a probabilistic information retrieval model widely used in the field of text retrieval.
-BM25 takes into account both term frequency and document length normalization, making it a robust and effective ranking model.
-(2) Dense retrieval methods:
-  • ANCE (Xiong et al., 2020) is a contrastive learning based model for ad-hoc dense retrieval.
-It selects hard training negatives from the entire corpus to reduce the stochastic gradient variance, and improve retrieval performance.
-  • BERT+ANN (Devlin et al., 2018; Chalkidis et al., 2020) is a strong baseline in ad-hoc retrieval tasks that uses BERT to encode documents and then apply Approximate Nearest Neighbor (ANN) search algorithms to retrieve relevant document.
-In this work, we adopt the BERT pre-trained on a large Chinese legal corpus [^7] https://github.com/thunlp/OpenCLaP..
-  • Lawformer+ANN (Xiao et al., 2021a): similar to BERT+ANN, we adopt Lawformer, which is a longformer (Beltagy et al., 2020) backbone and pre-trained on large legal case corpus, to encode legal cases.
-  • SAILER (Li et al., 2023) is a structure-aware pre-training method for legal case retrieval.
-It adopts an asymmetric encoder-decoder architecture to integrate structures of legal case document information into dense vectors.
-SAILER is the state-of-the-art legal case retrieval method.
-(3) Generative retrieval methods:
-  • DSI (Tay et al., 2022) is a new paradigm for document retrieval tasks.
-It utilizes a Transformer-based encoder-decoder model to map queries directly to relevant identifiers.
-DSI achieves the end-to-end retrieval.
-  • NCI (Wang et al., 2022) improves DSI in terms of using constrained beam search and prefix-aware weight-adaptive decoder.
-  • DSI-QG (Zhuang et al., 2022) design a query generation process on the top of DSI, which can mitigate data distribution mismatches present between the indexing and the retrieval phases.
-  • Ultron (Zhou et al., 2022) improves DSI through adopting product quantization to create semantic-based identifier and using URL to create term-based identifier.
-A.3. Implementation Details
-We implement baseline methods following the suggestions in the original papers.
-Specifically, (1) For classical term-based baselines, we use the pyserini [^8] https://github.com/castorini/pyserini.
-and genism [^9] https://github.com/RaRe-Technologies/gensim.
-toolkits with the default parameters.
-(2) For dense retrieval baselines, we use the open-sourced checkpoint [^10] https://github.com/thunlp.
-to initialize model parameters of BERT and Lawformer and use faiss [^11] https://github.com/facebookresearch/faiss.toolkit to implement ANN algorithms.
-The batch size is set to 16, the max length of input case is set to 512 for BERT and 1024 for Lawformer.
-We tune these models with in-batch contrastive loss.
-(3) For generative retrieval baselines, we directly use their official open-source implementations with the Chinese pre-trained T5 “Randeng”  [^12] https://huggingface.co/IDEA-CCNL/Randeng-T5-77M-MultiTask-Chinese.
-as backbone.
-We use beam search to retrieve relevant cases, where the beam size is set to 30.
-We keep the backbone model and beam size same as baselines and set the max input length of Gear to 512, the rest hyperparameters are tuned as follows: the batch size is set to 2; the learning rate is tuned from [1e-5,1e-4] with step size 2e-5; in the rationale extraction module, $k_{1}$, $k_{2}$, $k_{3}$ are respectively set to 2, 5, 15 for ELAM and 10, 20, 15 for LeCaRDv2 ; $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ are respectively set to 10.0, 1.0, 0.1; in the law structure constraint tree module, the height of the tree (the length of the hierarchical identifier) $L$ is set to 4; in the training, $\mu$ is set to 1; $\gamma_{h}$ is tuned from $\{0.01,0.1\}$; $\gamma_{h}$ is set to 1; $\lambda_{w}$ is tuned from $\{1,10,100\}$; $\lambda_{l}$ is tuned from [1e-4,1e-2] with step size 5e-4.
-Hyperparameters of Gear are tuned using grid search with Adam(Kingma and Ba, 2014).
-ALL experiments are conducted on a single Nvidia A6000.
-Figure 7.
-Comparison between the rationale extracted by Gear and the queries generated by Direct indexing, Doc2query, and ChatGPT.
-Underlined snippets denotes the key circumstances or key elements for the case.
-Figure 8. A case study of prompting ChatGPT to obatin the key elements and key circumstances for a charge.
-Appendix B Supplementary Experiments
-B.1. Case Study in Rationale Extraction
-In this section, we employ a case study to compare the performance of Gear and baseline methods in rationale extraction (query generation).
-Figure 7 showcases the comparison with all results translated from Chinese.
-We can find that (1) queries generated by Direct indexing and Doc2query fail to articulate the key circumstances and key elements of the original case, which are crucial factors for evaluating the relevance in legal case retrieval.
-This is because they lack law knowledge and struggle to handle the extremely lengthy documents.
-(2) The quality of both rationales extraction and ChatGPT summary is considerable as the underlined snippets highlighted.
-While both convey a substantial amount of key information, the rationales extracted by Gear are more sufficient and organized, which is easier for model learning.
-It contributes to our multi-level extraction strategy where we respectively extract keywords and sentences according to the definition of laws.
-The above results validate the effectiveness of the rationale extraction of Gear.
-B.2. Building Corpus by Prompting ChatGPT
-To improve the effectiveness of rationale extraction, we prompt chatGPT and extend the corpus using chatGPT’s feedback.
-Our prompt is shown in Figure 8.
-We would like ChatGPT play the role of a legal expert targeting on the Criminal Law and ask ChatGPT output the key circumstances and key elements for each charge.
-From the case, we found that chatGPT can provide multiple perspectives on a charge, including perpetrator, victim, behavior, demands, etc.
-These annotations serve as supplementary clarifications to the Criminal Law definitions and help Gear accurately select the informative rationales from legal cases.
+In this study, we introduce GEAR, a novel law-guided generative legal document retrieval method that explicitly integrates judgment prediction.
+GEAR exploits the law knowledge and extracts rationales from legal documents, ensuring a shared and informative representation for both tasks.
+Grounded in the inherent hierarchy of laws, GEAR constructs a law structure constraint tree and assigns the law-aware semantic ID to each document.
+These designs enable a unified traversal from the root, through intermediate charge nodes, to case-specific leaf nodes, which empowers GEAR to perform dual predictions for judgment and relevant documents in a single inference.
+With the help of the proposed revision loss, GEAR jointly minimizes the discrepancy between the IDs of predicted and labeled judgments/ retrieved documents, improving the accuracy and consistency for both tasks.
+Extensive experiments on two LCR datasets show the superiority of GEAR over state-of-the-art methods while maintaining competitive judgment prediction performance.
+Moreover, we validate its robustness across languages and domains on a French SAR dataset.
+Acknowledgements.
+This work was funded by the National Key R&D Program of China (2023YFA1008704), the National Natural Science Foundation of China (No.
+62376275, No.
+62377044), Beijing Key Laboratory of Big Data Management and Analysis Methods, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, PCC@RUC, funds for building world-class universities (disciplines) of Renmin University of China.